### v1

# How Much Content Do LLMs Generate That Induces Cognitive Bias in Users?

Abstract
Large language models (LLMs) are increasingly integrated into applications ranging from review summarization to medical diagnosis support, where they affect human decisions. Even though LLMs perform well in many tasks, they may also inherit societal or cognitive biases, which can inadvertently transfer to humans. We investigate when and how LLMs expose users to biased content and quantify its severity. Specifically, we assess three LLM families in summarization and news fact-checking tasks, evaluating how much LLMs stay consistent with their context and/or hallucinate. Our findings show that LLMs expose users to content that changes the sentiment of the context in 21.86% of the cases, hallucinates on post-knowledge-cutoff data questions in 57.33% of the cases, and primacy bias in 5.94% of the cases. We evaluate 18 distinct mitigation methods across three LLM families and find that targeted interventions can be effective. Given the prevalent use of LLMs in high-stakes domains, such as healthcare or legal analysis, our results highlight the need for robust technical safeguards and for developing user-centered interventions that address LLM limitations.
How Much Content Do LLMs Generate That Induces Cognitive Bias in Users?
Abeer Alessa, Akshaya Lakshminarasimhan, Param Somane, Julian Skirzynski, Julian McAuley, Jessica Echterhoff UC San Diego {aalessa, athenkarailakshminar, psomane, jskirzynski, jmcauley, jechterh}@ucsd.edu
1 Introduction
LLMs perform well across numerous tasks Albrecht et al. (2022) like content summarization Laban et al. (2023) , translation Elshin et al. (2024) , question-answering Lin et al. (2025), and sentiment analysis Zhang et al. (2024) including tasks with extremely large context Beltagy et al. (2020); Press et al. (2021); Touvron et al. (2023); Chen et al. (2023); Peng et al. (2023). In many of those tasks, humans rely on LLMs for daily decision-making support Rastogi et al. (2023); Li et al. (2022) as well as in expert contexts such as for writing policy documentsChoi et al. (2024) or summarizing medical documents (Spotnitz et al., 2024). However, these models have been shown to exhibit several societal biases Zhao et al. (2018); Nadeem et al. (2020); Liang et al. (2021); He et al. (2021), e.g., models favor specific genders Zhao et al. (2018), ethnicities Caliskan et al. (2017), or fall into similar decision-making-patterns as cognitively biased humans Echterhoff et al. (2024). Our specific contributions are the following
Figure 1: Our work evaluates how LLMs alter the source text, e.g., by changing its sentiment, to subsequently transfer the introduced bias to humans and affect their decision-making. Specifically, we evaluate how LLMs filter source content which leads to exposure bias for the user, how LLMs reframe the text which lead to human framing bias and how LLMs hallucinate which leads to authority/confirmation bias.
Figure 2: When LLMs process content for users to consume, they may change this content, e.g., by changing its sentiment or omitting some relevant parts. Exposure to the altered content may elicit cognitive biases such as positive framing bias or primacy bias, and subsequently have humans make different decisions than they would take if they saw the original text.
  1. 1.
    We analyze the character and extent of content alteration LLMs introduce in summarization and question-answering tasks. We find that LLMs change the framing of the text in roughly one of every five summaries and overly focus on the beginning of a source text once every 20 summaries. We also find that LLMs hallucinate in question-answering tasks in over half of the cases.
  2. 2.
    We propose specific metrics to capture the emergence of these content changes, such as a primacy bias score that captures the relative focus of the summary at the beginning of the source.
  3. 3.
    We provide a dataset of almost 3000 instances to analyze model hallucination past a model’s knowledge cutoff.  [^1] We publish this data at https://huggingface.co/datasets/sparklessszzz/NewsLensSync.
  4. 4.
    We evaluate 18 distinct mitigation methods and discover effective methods that target each specific type of content alteration.
Our goal is to highlight LLMs’ unreliability in various tasks and stress the importance of mitigating their errors, as they can negatively influence human users. We hope that our work will serve as a stepping stone for creating methods that improve LLMs and interventions that teach humans how to collaborate with LLMs thoughtfully.
2 Related Work
2.1 Social and Representational Biases in Language Models
Previous work shows the existence of different biases in LLMs. For example, prior work explores ways to evaluate human-like cognitive biases Jones and Steinhardt (2022); Echterhoff et al. (2024). Primacy and recency bias in LLMs have been shown in question-answering settings, where the models tend to be more accurate when the answer is presented in earlier or later documents Liu et al. (2024); Peysakhovich and Lerer (2023); Wang et al. (2023); Zhang et al. (2023); Eicher and Irgolič (2024). When LLMs are presented with options (e.g., A/B/C/D), they are shown to be more sensitive to the order Zheng et al. (2023a). Other studies have demonstrated the existence of framing bias in LLMs Leto et al. (2024). In framing bias, alterations in the tone, e.g., from positive to negative, can change the model’s decision-making Jones and Steinhardt (2022); Echterhoff et al. (2024).
In contrast to prior work that evaluates biases inherent in model behavior, we quantitatively assess the extent to which LLM responses alter the underlying content and how much bias they emit through these modifications.
2.2 Cognitive Effects of LLM-Generated Content on Humans
LLMs generate content to help users with all kinds of different tasks, such as summarization Laban et al. (2023), coding Tong and Zhang (2024), and writing Spangher et al. (2024) for personal or professional Draxler et al. (2023) use. When a model is given a specific task, it can alter, omit, or add information to produce the final model output. This can have an impact on a user in several ways. Previous work shows that LLMs can influence users by reinforcing users’ existing opinions Sharma et al. (2024), which can increase users’ tendency towards confirmation bias. The style in which LLMs present advice can significantly influence user perceptions Wester et al. (2024).
Prior work shows that altered headlines can introduce a framing effect on people in collective decision-making Abels et al. (2024). When experts use LLMs, they accept their suggestions without much revision Choi et al. (2024)
Given this insight, our study evaluates framing changes that the model introduces when generating a summary from ground-truth content, which subsequently could bias a user.
3 Background
LLMs may amplify different cognitive biases when users interpret their outputs. We focus on four biases:
Framing bias refers to changes in human decisions when seeing a problem presented differently. People make different choices based on the wording of the questions, even if the actual choices presented remain unchanged Tversky and Kahneman (1981). Based on this prior work, we assume that a change in framing impacts the user and is hence to be avoided, e.g., when comparing the framing of a ground-truth context and its summarization.
Primacy and Recency Bias. In primacy bias, humans prioritize the information they encounter first, while in recency bias, humans prioritize information they encountered most recently Murphy et al. (2006). This can result in skewed decision-making regardless of relevance or accuracy, but simply by order of encountered information Glenberg et al. (1980). When LLMs fall prey to primacy/recency bias Echterhoff et al. (2024), they might filter or feature information to the user leading to exposure bias. We assume that systematically highlighting any part of the context, whether it appears first, middle, or last, can influence the human user and should therefore be avoided.
Confirmation and Authority Bias. Humans favor information confirming their beliefs or expectations Nickerson (1998). We assume that when a user prompts the model with belief-consistent input, and the model hallucinates in response, this can reinforce the user’s prior beliefs and lead to biased or misinformed decisions. In authority bias, people are more likely to trust information from perceived experts or authoritative sources Milgram (1963); Cialdini (2007); Güner Gültekin (2024). Research on human-AI interaction suggests that similar dynamics apply when people interact with artificial systems perceived as experts, which can create challenges when verifying facts Glickman and Sharot (2025). For example, when a user asks an LLM to verify if/when a particular legislation was passed, a user with strong prior beliefs may exhibit confirmation bias, while others may default to authority bias.
4 Quantifying Content Alteration in LLM Responses
Due to the effects of the aforementioned biases, we propose the following metrics to capture unintended content alterations introduced by LLMs.
4.1 Framing Effects in Model-Generated Content
We test if a model changes the sentiment of its summary compared to the context. For example, a model shifts the summary sentiment to positive, whereas the ground truth context was neutral. To evaluate these changes, we classify the framing of context $f_{c}$ (positive, negative, or neutral) before a summarization task, as well as on the generated answer summary by the model $f_{s}$. We use GPT-4o-mini to classify the framing of a text and provide details of its accuracy in Appendix A. We consider a model to be consistent if $f_{c}$ equals $f_{s}$. We consider the Bias-Inducing (BI) score as the Framing-change fraction, and denote it by $\varphi_{\mathrm{frame}}$ as below
$\varphi_{\mathrm{frame}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl{(}f_{c}^{(i)}\neq f_{s}^{(i)}\bigr{)},$ (1)
where $\mathbb{I}(\cdot)$ is the indicator function that returns one if the condition is true and zero otherwise.
4.2 Primacy Effects in Model-Generated Content
We assess whether the model emphasizes a particular part of the review. For example, a model focuses only on the beginning of an Amazon user review that begins with praise but ends with criticism. We split the original context into three equal parts and calculate the cosine similarity using the multilingual-E5 embedding Wang et al. (2024) between each part and the summary. We denote the similarities of the summary and the beginning as $s_{b_{i}}$, the middle as $s_{m_{i}}$, and the end as $s_{e_{i}}$ for each instance $i$. We define biased examples as those in which the similarity to the beginning exceeds the similarity to the middle by a threshold of at least $\alpha$. We define the Primacy Bias Score as:
$\psi_{\mathrm{pri}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(s_{b_{i}}>s_{m_{i}}+\alpha\right)$ (2)
with $\alpha=0.05$ for this study. A higher score indicates that a greater proportion of summaries focus more on the beginning of the source content compared to the middle.
4.3 Hallucination Effects in Model-Generated Content
We quantify whether a model generates hallucinated content when asked to fact-check news that lies outside its knowledge cutoff. For example, a model may wrongly state that an event never occurred, even though it did, but only after the model’s training had been completed. We use pre- and post-knowledge-cutoff data and prompt a model to verify whether specific news is true or false. We test both the accuracy of true, real-world news (Actual News Accuracy) and its falsified (negated) counterparts (Falsified News Accuracy). We prompt the model to evaluate the factual accuracy of a given news description. We propose to evaluate the strict accuracy in identifying when both the true news and the falsified news were correctly predicted.
For strict accuracy, we evaluate paired instances of true news and its falsified counterparts:
$s_{H}^{T}(D_{j})=\frac{1}{|P_{j}|}\sum_{(x_{i}^{t},x_{i}^{f})\in P_{j}}\mathbb{I}(\hat{y}_{i}^{t}\land\neg\hat{y}_{i}^{f})$ (3)
where $P_{j}$ is the set of paired instances in dataset $D_{j}$, $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{f}$ are the model’s predictions for the true and falsified versions of the same news respectively, and $T$ is the time-horizon (pre- or post-model-knowledge-cutoff).
We analyze the disparity between pre-cutoff and post-cutoff performance:
[…]
A larger $\delta_{H}$ indicates a greater chance of generating incorrect information when responding to queries beyond the model’s knowledge cutoff.
5 Mitigating Bias-Inducing Patterns in LLM Responses
To mitigate biased patterns in LLM responses, we build on our quantification of content alterations introduced in Section 4. We introduce targeted interventions designed to ensure balanced content coverage, reduce misleading emphasis, and maintain accurate representation in model outputs.
5.1 General Purpose Mitigations
Self-Awareness Prompt (SA).
Prior works suggest that simple awareness prompts can mitigate certain biases Mair et al. (2014); Echterhoff et al. (2024). For example, in the summarization setting, we use the following prompt:
“You are an unbiased summarizer; be mindful not to omit coverage of any portion of the text or alter its sentiment. Do not incorporate framing bias.”
Chain-of-Thought Prompting (CoT).
Chain-of-Thought prompting encourages the model to decompose complex problems by generating a series of intermediate reasoning steps before producing a final answer Wei et al. (2023). In the context of summarization, we guide the model to articulate steps for processing different parts of the input text. For the question-answering setting, we explicitly define when to return specific responses during fact-checking.
5.2 Primacy Effects Mitigation
We group interventions into four families: Prompt-level cues, chunk-based summarization, attention/order re-ranking, and decoding-time control. Each family targets the tendency of LLMs to over-weight early context (primacy) or flip the source sentiment. A compact reference of all the methods that we experimented with is given in Appendix C.
Stand-out methods.
Weighted Summaries is a chunk-based approach that ensures balanced representation from the source text. The algorithm first splits the document into three segments (beginning, middle, and end). It then generates a partial summary for each segment independently, constraining the output length for each part according to a pre-allocated token budget (33%, 34%, and 33% of the total desired summary length). These partial summaries are then combined to form the final output.
In the Mirostat Decoding approach, the model’s sampling randomness (temperature) is dynamically adjusted during generation. This ensures the output text maintains a consistent level of unpredictability, which can prevent the model from focusing too narrowly on initial sentences and encourage broader coverage of the source text. We use a target entropy of $\mu^{\star}\!=\!2.0$ in our runs.
5.3 Mitigation for Framing Effects
In the Weighted Token Decoding approach, we directly influence the model’s word choice during generation. At each step, we modify the output probabilities by adding a logarithmic weight to the logits of specific tokens before sampling, in particular we down-weight tokens associated with negative sentiment to reduce framing shifts, as detailed in Appendix C. Liu et al. (2021); Dathathri et al. (2020)
5.4 Mitigation for Hallucination
The mitigation strategies for hallucination are summarized in Appendix C.
Knowledge Boundary Awareness. We propose a novel approach that explicitly defines the model’s knowledge cutoff date, preventing the fabrication of information about events beyond training data.
Epistemic Tagging. This approach requires models to express confidence levels alongside factual assertions, which forces self-evaluation of fact-checking abilities (i.e., the model additionally respond with its confidence level high or low for each of its responses.) Lin et al. (2022).
All methods are prompt-based interventions. Prompts details are in Appendix E.
Amazon Reviews MediaSum News
Method $\varphi_{\mathrm{frame}}$downarrow $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\varphi_{\mathrm{frame}}$downarrow $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$
GPT-3.5-turbo 16.0% 0.848 0.826 0.825 7.6% 24.8% 0.840 0.823 0.820 6.1%
Llama-3.2-3B-Instruct 14.9% 0.860 0.842 0.840 7.4% 22.1% 0.851 0.837 0.832 5.1%
Llama-3-8B-Instruct 14.5% 0.855 0.837 0.836 7.0% 21.9% 0.847 0.834 0.828 4.0%
Phi-3-mini-4k-Instruct 34.5% 0.836 0.822 0.823 5.4% 26.2% 0.828 0.827 0.823 4.9%
Table 1: Models introduce framing changes ($\varphi_{\mathrm{frame}}$) and filter/feature content based on position ($\psi_{\mathrm{pri}}$). $\varphi_{\mathrm{frame}}$ is the proportion of examples where the framing changed from the original to the summary (lower is better). $\bar{s}_{b}$, $\bar{s}_{m}$, and $\bar{s}_{e}$ represent the average cosine similarity between the summary and the beginning, middle, and end thirds of the source text, respectively. Higher values of $\bar{s}_{b}$, $\bar{s}_{m}$, and $\bar{s}_{e}$ indicate stronger content preservation from those respective sections. $\psi_{\mathrm{pri}}$ represents the percentage of examples with primacy bias as we defined in Equation 2.
6 Experiments
6.1 Datasets
News Interviews
News summaries should have broad coverage of the content, independent of content position, as all of what has been said is potentially relevant for a summary. As these interviews are supposedly an objective depiction of current states and affairs, a summary should not be more positive or negative than the original content. We use MediaSum, a media interview dataset for dialogue summarization from NPR and CNN Zhu et al. (2021) to measure framing and primacy effects. We randomly select 1000 data points with tokens $\leq 4000$ to have the full content fit into a model’s context.
Product Reviews
Previous work suggests that changing the framing of the product influences consumer purchase decisions Wei et al. (2024). To explore this effect in the context of LLMs outputs, we use the Amazon Reviews dataset Ni et al. (2019). It contains customer reviews from Amazon across various product categories, providing information on product quality, usability, and satisfaction Ni et al. (2019). We sample 1,000 random examples from the Electronics category, filtering for token lengths $\leq 4000$ to maintain full-context input for the model.
News Hallucination
We provide a dataset to quantify how models can discern factual information across different temporal boundaries and in the true and negated versions of a news piece.
Pre-Knowledge Cutoff Data. For the pre-knowledge cutoff evaluation, we leverage the “All the news” data all (2020), which includes articles collected from 27 major American publications from 2016 to 2020. This data is available for non-commercial research purposes. We sampled 2,700 random data points in our experiment.
Post-Knowledge Cutoff Data. To evaluate model performance on recent events (i.e., events not included in a model’s training data), we introduce a self-updating dataset, the NewsLensSync Akshaya (2025), specifically designed to measure ongoing hallucinations for content beyond models’ knowledge cutoff dates. We evaluate 2801 real news descriptions and their negated versions.
For both post- and pre-cutoff examples, we include the true news events and carefully crafted counterfactual versions. We used a transformer-based model to generate the counterfactual examples Anschütz et al. (2023).
6.2 Models
We use four open- and closed-source language models. Details on the models and all of our hyperparameter settings can be found in Appendix B.
7 Results
7.1 LLMs Alter the Framing of Reviews when Summarizing
[…]
MediaSum GPT-3.5-turbo 2.9% 8.6% 0.1% 0.1% 1.9% 11.2%
Llama-3-8B-Instruct 2.6% 8.0% 0.0% 0.1% 2.3% 8.9%
Llama-3.2-3B-Instruct 2.0% 7.7% 0.1% 0.1% 2.3% 9.9%
Phi-3-mini-4k-Instruct 2.6% 7.3% 0.1% 0.7% 2.4% 13.1%
Amazon GPT-3.5-turbo 2.6% 0.5% 0.2% 1.4% 7.4% 3.9%
Llama-3-8B-Instruct 2.3% 0.6% 0.7% 1.4% 6.7% 2.8%
Llama-3.2-3B-Instruct 1.8% 0.6% 0.5% 2.1% 6.2% 3.7%
Phi-3-mini-4k-Instruct 1.6% 0.8% 1.8% 2.2% 21.2% 6.9%
[…]
We observe that models change the text framing of their summary compared to the context in both media interviews and the Amazon dataset (Table 1). Most models (Llama and GPT-3.5-turbo) show a framing change fraction between 14.5% and 16%, whereas Phi-3-mini-4k-Instruct shows 34.5% framing changes. In Table 3, we show examples of a framing change between a summary and the ground-truth context. Table  2 includes a detailed evaluation of framing alterations (e.g., from neutral to positive, or positive to negative). In the MediaSum news interviews, most framing shifts are from neutral to negative between 7.3% and 8.6% and negative to neutral between 8.9% and 13.1%. In Amazon reviews, the most common shift is from positive to neutral between 6.2% and 7.4% for Llama and GPT-3.5-turbo models and 21.2% for Phi-3-mini-4k-Instruct. This suggests that the models tend to downplay positive sentiment in product reviews. This pattern is consistent across all evaluated models.
Original Review GPT-3.5-turbo Summary
“This is a fun, useful tablet for the price. I love that you can make all different homepages. Love that you can pick wallpaper which you can’t do with Kindle Fires, which are the only other tablets I’ve owned. Of course, if you want a tablet mainly for books, Kindle is the tablet you want. I also love all the apps you can get with the Samsung. What I don’t like is the 8 GB storage. I was under the impression that if you got the 32 GB extra, it would integrate with the tablet rather than just be like a drive for MP3s and such. 8 gigs is barely enough for the apps you need. Also, the manual tells you nothing about how to add things to the extra drive. Had to go to YouTube. My tablet has connection issues from time to time. I still really enjoy it most of the time. It can be super fast and plays things well from the cloud. Very light and attractive. I am returning the tablet. 8 GB is just not enough. Extra storage is cumbersome to use and doesn’t work half of the time. Not easy like the cloud. Speakers unimpressive. I am disappointed because I thought I would like the Samsung. Also a battery hog. Will try the Nexus 7.” “I enjoyed the Samsung tablet for its customization options and app variety, but the 8 GB storage was insufficient. The extra storage was cumbersome and unreliable. Connection issues and battery drain were also drawbacks. I will be returning it and trying the Nexus 7 instead.”
Table 3: The framing of this example was classified as neutral before the summarization task to negative after.
7.2 LLMs Summarization have Imbalanced Context Coverage
We find that on average, summaries exhibit greater alignment with the beginning of the source text compared to the middle and end sections (Table 1). For example, across all models and datasets, the average similarity to the beginning section ranges from 0.828 to 0.860, which is consistently higher than the similarity to the middle 0.822-0.842 and end 0.820-0.840 sections. This coverage imbalance is further quantified using our primacy bias score, which captures the proportion of examples where beginning content is prioritized by at least 5% over other middle section. For the Amazon Reviews dataset, $\psi_{\mathrm{pri}}$ ranges from 5.4% to 7.6%. Similarly, MediaSum summaries exhibit $\psi_{\mathrm{pri}}$ scores between 4.0% and 6.1%, which shows primacy bias even in a more structured news content.
7.3 LLMs Hallucinate Post Knowledge-Cutoff
Baseline results in Table 4 show consistent drops in strict accuracy $(s_{H}^{T})$ from pre-cutoff to post-cutoff data across all models. Llama-3-8B-Instruct declined from 26% to 21% ($\delta_{H}=5\%$), Llama-3.2-3B-Instruct dropped from 19% to 13% ($\delta_{H}=6\%$), and Phi-3-mini-4k-Instruct showed a $\delta_{H}=4\%$ decline.
Models frequently hallucinate or misclassify new events, highlighting their inability to distinguish true from false statements without recent training data. Llama models averaged a $\delta_{H}=5\%$ decline despite individual accuracies above 40% on both real and falsified news. Llama-3-8B-Instruct performed best on true events across both settings, while Phi-3-mini-4k-Instruct showed poor performance with only 8% strict accuracy $(s_{H}^{T})$ on real news data.
Models Prompt Strategy Pre-cutoff data Post-cutoff data
Actual News Acc. Falsified News Acc. Strict Acc. $(s_{H}^{T})$ Actual News Acc. Falsified News Acc. Strict Acc. $(s_{H}^{T})$
Llama-3-8B-Instruct Baseline 74% 46% 26% 75% 42% 21%
Prompt Calibration (CoT) 39% 85% 27% 36% 84% 22%
Knowledge Boundary Aware 65% 54% 25% 50% 63% 18%
Epistemic Tagging 80% 35% 18% 78% 37% 19%
Llama-3.2-3B-Instruct Baseline 48% 58% 19% 35% 72% 15%
Prompt Calibration (CoT) 48% 56% 17% 46% 63% 15%
Knowledge Boundary Aware 25% 87% 15% 15% 94% 10%
Epistemic Tagging 51% 68% 26% 53% 64% 23%
Phi-3-mini-4k-Instruct Baseline 14% 97% 12% 9% 98% 8%
Prompt Calibration (CoT) 3% 99% 3% 3% 99% 3%
Knowledge Boundary Aware 3% 99% 3% 2% 99% 2%
Epistemic Tagging 45% 76% 31% 53% 70% 29%
Table 4: Accuracies on real-news items and their falsified counterparts, evaluated before and after each model’s training cutoff date. Strict Accuracy $s_{H}^{T}(D_{j})$ counts only predictions where the model gets both members of each (True, False) pair correct.
7.4 Special Biases Require Special Mitigations
The 18 mitigations we study (definitions in Appendix C) do not alleviate all biases simultaneously. Each method involves trade-offs among all biases, and its effectiveness varies depending on the model and corpus. For example, techniques aimed at reducing position-based salience (e.g., primacy bias) often work by redistributing attention across the input. However, this can disrupt how the model maintains consistent sentiment in its responses.
Method $\bar{s}$ $\psi_{\mathrm{pri}}$$\varphi_{\mathrm{frame}}$§
Baseline 0.843 7.0% 14.5%
Weighted Summaries 0.910 15.1% 15.1%
Mirostat Decoding 0.902 15.5% 13.8%
Pos.-Invariant Shuffle 0.891 11.5% 21.2%
Weighted Token Decoding 0.858 18.4% 13.6%
Self-Awareness Prompt 0.889 22.9% 15.2%
Table 5: Key mitigation results on Amazon Reviews for Llama-3-8B-Instruct. Mean content-coverage similarity $(\bar{s}_{b}+\bar{s}_{m}+\bar{s}_{e})/3$. Primacy-bias score. §Framing-change rate. Lower is better for $\psi_{\mathrm{pri}}$ and $\varphi_{\mathrm{frame}}$. Behavior on other corpora/models follows a similar qualitative pattern; see full results in Tab. 9.
7.4.1 Primacy Bias and Coverage
The Weighted Summaries method, which assigns a fixed token budget to each text section, consistently increases overall content coverage. For instance, with Llama-3-8B-Instruct on Amazon Reviews, it boosts average coverage similarity from a baseline of 0.843 to 0.910. However, this comes at the cost of worsening the primacy bias score ($\psi_{\mathrm{pri}}$), which increases from 7.0% to 15.1%.
Mirostat Decoding also lifts coverage; it is the only method that cuts $\psi_{\mathrm{pri}}$ on the small Phi-3 model (-1.2 pp) but worsens it on larger models.
Position-Invariant Shuffle is useful diagnostically because it helps confirm that the model relies on word order. By shuffling sentences randomly, we remove positional cues; the resulting change in output shows the model’s sensitivity to input structure. While this method increases overall coverage (e.g., from 0.843 to 0.891 for Llama-3-8B), it degrades other metrics, increasing the primacy score $\psi_{\mathrm{pri}}$ to 11.5% and the framing change rate $\varphi_{\mathrm{frame}}$ to 21.2%.
7.4.2 Framing Consistency
Weighted Token Decoding proves most effective for this bias. By down-weighting negative lexemes during decoding, it delivers the lowest framing-change fraction ($\varphi_{\mathrm{frame}}$) for Llama-3-8B, reducing it by 0.9 percentage points to 13.6%. However, this targeted intervention significantly worsens primacy bias, with $\psi_{\mathrm{pri}}$ increasing from 7.0% to 18.4%.
Prompt-only nudges. Lightweight prompts such as Self-Awareness or Chain-of-Thought alter $\varphi_{\mathrm{frame}}$ by only a few percentage points in most settings. However, the Self-Awareness variant gives the best framing result on MediaSum for Llama-3-8B-Instruct, reducing $\varphi_{\mathrm{frame}}$ to 21.0% from a 21.9% baseline at a negligible computational cost.
7.4.3 Hallucination
Prompt Calibration using Chain-of-Thought. This mitigation strategy improved the strict accuracies for Llama-3-8B-Instruct by 1% on both pre and post cut-off data.
For Llama-3.2-3B-Instruct, the strict accuracy dropped by 2% on pre cutoff data and there was no significant improvement when applied on post cut-off data.
Phi-3-mini-4k-Instruct showed the most dramatic degradation. Real news accuracy plummeted from 14% to 3% on pre-cutoff data. Post-cutoff data showed similar degradation in real news detection while maintaining high falsified news detection.
The approach appears to work reasonably well for larger models like Llama-3-8B-Instruct but causes significant performance degradation in smaller models like Phi-3-mini-4k-Instruct, consistent with known limitations in factual recall for smaller language models Abdin et al. (2024).
Knowledge Boundary Awareness. Making Llama-3-8B-Instruct aware of its knowledge boundaries showed some significant improvements by 15% in falsified news detection and maintained it’s strict accuracy close to the baseline in pre-cutoff data. In case of Llama-3.2-3B-Instruct and Phi-3-mini-4k-Instruct models, the performance declined in pre- and post-cutoff date data.
Epistemic Tagging. This method performed the best across all models. It associates confidence level during the fact checking mechanism (i.e., highly or less confident) alongside answering whether an event has occurred or not (Table 6)
Phi-3-mini-4k-Instruct was able to outperform with a steep increase in the strict accuracy by 21%. The Llama-3.2-3B-Instruct also performed well with strict accuracies improved by 8% and a balanced independent accuracies for real and falsified news data above 50%. On the other hand, Llama-3-8B-Instruct’s accuracy in identifying real news improved and maintained the strict accuracy around the baseline.
Models cutoff Actual News Falsified News
High Low High Low
Llama-3-8B-Instruct Pre 99.0% 1.0% 99.8% 0.2%
Post 98.3% 1.7% 99.9% 0.1%
Llama-3.2-3B-Instruct Pre 75.8% 24.2% 24.6% 75.4%
Post 66.7% 33.3% 83.3% 16.7%
Phi-3-mini-4k-Instruct Pre 93.8% 6.2% 86.2% 13.8%
Post 98.1% 1.9% 95.4% 4.6%
Table 6: Epistemic Tagging prediction Confidence levels (high, low) for different models across cutoff conditions.
8 Conclusion
When LLMs modify the framing or emphasis of the content they process, they can inadvertently shape human perception and decision-making. In this work, we quantify how much LLMs alter original content in a summarization and news fact-checking task. We find that models, on average, introduce framing bias to the user in 21.86% of instances, measured by sentiment shifts between the source text and the generated summary and primacy bias in 5.94% of the cases, where summaries disproportionately reflect the beginning of the original text. The models also, on average, hallucinate on post-knowledge-cutoff data questions in 57.33% of instances.
Across the mitigation strategies evaluated, the effectiveness of each approach depends on the model type and the bias targeted. Weighted Summaries and Mirostat decoding showed promise in reducing framing changes and positional biases, particularly in smaller and mid-sized models. Epistemic Tagging consistently improved factual reliability on post-knowledge-cutoff knowledge.
This paper takes a step toward a careful analysis and mitigation of content alteration to avoid the risk of LLMs introducing systemic biases into decision-making processes across various domains, including media, education, and public policy.
9 Limitations
This study examines the extent to which LLM outputs alter original content in ways that may influence human judgment. While we ground our assumptions in prior cognitive science research, as discussed in the paper, we aim to validate these effects through a dedicated user study measuring how such alterations impact human decision-making.
Most of our mitigation methods were designed with one specific issue in mind (e.g. framing, hallucinations, or primacy effects). In our future work, we aim to explore more general mitigation techniques that address overlapping biases simultaneously.
Our work aims to encourage the NLP community to continuously test LLMs for biases that can influence users’ decision-making, which is why we introduce the self-updating NewsLensSync dataset Akshaya (2025). Currently the dataset holds articles related to "political" domain. Soon we will try to incorporate other areas to make it more generalizable and usable across diverse set of researchers. However, our data could be misused for training models to reinforce existing ideological narratives or use falsified data. We urge researchers to be cautious in framing and contextualizing their use of the data. We provide our data under the CC-BY-4.0 license.
Ethical Considerations.
Our analysis shows that LLMs pay more attention to content from the beginning of the context, hallucinate facts, and alter the text framing in their outputs. These alterations can have an impact on the user, and skew their decision-making. These findings underscore the risks of working with LLMs, even in everyday tasks.
Experiments. All experiments are run on NVIDIA RTX A6000 for open-source models and the official APIs for closed-source models with a fixed random seed. Our run time was approximately 360 GPU hours. All prompts details are included in Appendix E.
Appendix A LLMs as a framing judge
Previous work has shown that high-capability LLMs like GPT-4 can achieve over 80% agreement with human experts on question answering tasks, the level of agreement observed between humans themselves (81%). GPT-4 reached 85% agreement with human judgments, suggesting that LLMs can serve as effective proxies for human preferences Zheng et al. (2023b). Thus, we test three high-capability models to classify the framing ( GPT-3.5-turbo, GPT-4-turbo, and GPT-4o-mini). We experimented on a sample of 500 randomly selected Amazon reviews, using the user-provided rating as ground truth to evaluate the accuracy of the models. We asked the models to rate the product based on the review, then we mapped ratings 1 and 2 as negative, 3 as neutral, and 4 and 5 as positive. The GPT-3.5-turbo, GPT-4-turbo, and GPT-4o-mini achieved accuracy of 0.89, 0.91, and 0.92, respectively. Based on this, we selected GPT-4o-mini to serve as the framing judge.
Appendix B Model setup and Hyperparameters
Our implementation utilizes the Hugging Face Transformers library using the official APIs. We employ Flash Attention 2 when available to enhance computational efficiency during inference. For consistency across experimental conditions, we maintain fixed generation parameters with temperature=0.01, do_sample=False, and max_new_tokens= 500 set to a constant with fixed random seed. Table 7 shows the models we evaluated, their context window, and the rationale behind choosing them.
Model Context Window Reason
GPT-3.5-turbo 16k A closed-source, highly capable model that is cost-effective OpenAI (2024).
Llama-3-8B-Instruct 8k Outperforms competing models on both per-category win rate and average per-category score Grattafiori et al. (2024).
Llama-3.2-3B-Instruct 8k Lightweight yet powerful; offers multilingual text-generation abilities Meta AI (2024).
Phi-3-mini-4k-Instruct 4k Matches high-performance closed-source models on several key benchmarks Abdin et al. (2024).
Table 7: Models used in our evaluation, along with their context-window sizes and selection rationale.
Appendix C Detailed Description of Mitigation Methods
Table 8 provides a concise, side-by-side reference for all 18 mitigation approaches evaluated in this work, including their intended bias target(s), operational definition, key hyper-parameters, and the original citation.
Family Method Operational definition Target bias Key hyper-parameters Citation
Prompt Self-Awareness Prompt Prepend an explicit directive that the model remain neutral, preserve sentiment, and cover all parts of the source text. General Mair et al. (2014); Echterhoff et al. (2024)
Chain-of-Thought Prompting Model generates intermediate reasoning steps for each segment before emitting a final summary, encouraging balanced coverage. General Wei et al. (2023)
Cloze-Style Prompt Insert tags BEGIN:__, MIDDLE:__, END:__; model fills blanks, then emits FINAL_SUMMARY. Primacy Tags inside prompt; no extra threshold. Liu and Lapata (2019)
Cognitive Counterfactual Simulation Draft $\rightarrow$ imagine how primacy, recency, or framing bias would distort it $\rightarrow$ rewrite to avoid those distortions. Primacy,   Framing Two passes. Dong et al. (2023)
Self-Help Debias Draft, model self-critiques positional coverage, rewrites summary. Primacy Rewrite pass 300 tokens. Echterhoff et al. (2024)
Task-Specific CoT Prompt Calibration Establishes clear evaluation criteria by explicitly specifying when to return “True” or “False” during fact-checking tasks. For example: “Return true only if the described event has occurred, or if it is a direct consequence of a previously known event”. Return false in all other cases Hallucination Chain-of-Thought sequence pertaining to your task Wei et al. (2023)
Knowledge Boundary Awareness Explicitly defines the model’s knowledge cutoff date to create clear temporal boundaries for what the model can reasonably be expected to know. Prevents models from fabricating information about events beyond their training data by acknowledging these boundaries in prompts. Hallucination Knowledge cutoff date specification for the model being employed
Epistemic Tagging Requires models to express confidence levels alongside factual assertions, creating a more nuanced representation of the model’s knowledge state. Forces the model to evaluate its own fact-checking abilities before responding and provides users with meta-cognitive signals about response reliability. Hallucination Confidence level scales - in this case we used High and Low. This can incude moderate, very high, very low depending on task. Lin et al. (2022)
Chunk Partial Summaries Ensemble Split the document into equal-length chunks and summarize each chunk independently (same generation hyperparameters). In a second pass, we merge the partial summaries into a final summary. Primacy Chunk size $\lfloor|D|/3\rfloor$; greedy merge. Ou and Lapata (2025)
Weighted Summaries We pre-allocate the summary length budget so that 33% focuses on the beginning, 34% on the middle, and 33% on the end. The model is instructed not to exceed these chunk limits, which helps prevent overemphasis on earlier parts. Primacy Token ratio $0.33{:}0.34{:}0.33$. inspired by Liu and Lapata (2019)
Re-rank Attention-Sort Re-ordering Run a forward pass (without generating the final summary) to estimate cross-attention weights for each paragraph or segment, then reorder segments from lowest-attended to highest-attended. The model is then re-prompted with this new order, which pushes under-attended parts later in the context and encourages more balanced coverage. Primacy 2 iterations; paragraph granularity.
Position-Invariant Shuffle Shuffle entire sentences (split by periods) to randomize their order before prompting the model, so it cannot rely on absolute position. This can disrupt semantic continuity slightly, but ensures that coverage does not depend on text location. Diagnostic ablation for positional sensitivity. Primacy Period-split; seed $42$. Wan et al. (2024); Liu et al. (2024)
Decode Forced Balanced Coverage We measure coverage of the beginning, middle, and end sections via TF-IDF Salton and Buckley (1988), and add $\log\gamma$ ($\gamma=1.5$) to logits of tokens from under-covered sections until $\lvert s_{b}\!-\!s_{e}\rvert\leq 0.05$. Primacy Threshold $0.05$; boost $\gamma=1.5$. See et al. (2017)
Weighted Token Decoding Multiply next-token probabilities by weights $w_{i}$ (down-weight negative words, up-weight middle keywords). Framing $w_{\text{neg}}=0.3$, $w_{\text{mid}}=2.0$. Liu et al. (2021); Dathathri et al. (2020)
Mirostat Decoding After each token, compute its surprise $s_{t}=-\log p_{t}$; update the running state with $\mu_{t+1}=\mu_{t}-\eta\,(s_{t}-\mu^{\star})$, set temperature $T_{t}=\exp(\mu_{t+1})$, and rescale logits as $\mathrm{softmax}(z/T_{t})$. The feedback loop keeps the observed surprise near the target $\mu^{\star}$, smoothing positional coverage. Primacy $\mu^{\star}=2.0,\;\eta=0.1$ Basu et al. (2021)
Rejection Sampling If top-1 token would increase chunk imbalance, set its logit to $\infty$- ∞ and resample within top-$k$. Primacy $k=5$.
Self-Debias Decoding Dual-pass decoding: at each step run a second forward pass on the current context preceded by a bias-inducing prefix that names the undesired attribute; obtain $p_{\text{bias}}$, compute $\Delta=p_{\text{main}}-p_{\text{bias}}$, and down-scale tokens with $\Delta<0$ via $\alpha(\Delta)=e^{\lambda\Delta}$. Framing $\lambda=10$; bias prefix $<\!30$ tokens; refresh every 4 steps. Schick et al. (2021)
Local-Explanation Guard After each token the model explains its choice; if explanation says “ignoring middle” or “flipping sentiment”, reject token and resample. Primacy,   Framing Check every 5 steps; rule-based detector.
Table 8: Overview of the 18 bias-mitigation strategies evaluated in this work. Methods are grouped by the family that they belong to, annotated with the bias they aim to address, and accompanied by the main hyper-parameters and primary citation(s) used in our experiments.
Appendix D Full Coverage, Framing, and Bias Metrics
Table 9 reports the complete results for all models and mitigation strategies.
Amazon Reviews MediaSum News
Method $\varphi_{\mathrm{frame}}\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\rho_{\text{pri}}\downarrow$ $\varphi_{\mathrm{frame}}\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\rho_{\text{pri}}\downarrow$
GPT-3.5-turbo
Baseline 16.0% 0.848 0.826 0.825 7.6% 24.8% 0.840 0.823 0.820 6.1%
Llama-3.2-3B-Instruct
Baseline 14.9% 0.860 0.842 0.840 7.4% 22.1% 0.851 0.837 0.832 5.1%
Self-Awareness Prompt 24.0% 0.908 0.883 0.881 19.7% 80.4% 20.9% 0.907 0.883 0.875 15.7% 85.0%
Chain-of-Thought 23.6% 0.909 0.883 0.881 19.9% 81.1% 22.6% 0.907 0.883 0.876 16.9% 85.9%
Cloze-Style Prompt 23.8% 0.908 0.883 0.881 19.8% 81.0% 22.6% 0.907 0.883 0.876 16.9% 85.9%
Cognitive Counterfactual Sim. 23.7% 0.908 0.883 0.881 20.2% 80.8% 22.5% 0.907 0.884 0.876 16.6% 84.8%
Self-Help Debias 23.0% 0.908 0.883 0.881 20.6% 81.3% 23.6% 0.908 0.884 0.876 15.9% 85.8%
Partial Summaries Ensemble 21.0% 0.886 0.870 0.869 14.8% 69.1% 27.6% 0.871 0.854 0.846 14.3% 78.2%
Weighted Summaries 20.7% 0.921 0.903 0.899 15.4% 75.6% 29.4% 0.876 0.862 0.852 9.8% 80.1%
Attention-Sort Re-ordering 19.8% 0.905 0.886 0.885 16.8% 71.4% 26.4% 0.859 0.841 0.830 14.6% 78.5%
Position-Invariant Shuffle 26.6% 0.903 0.890 0.889 10.8% 67.1% 30.8% 0.875 0.868 0.859 8.9% 68.4%
Forced Balanced Coverage 21.0% 0.905 0.884 0.883 17.5% 72.5% 26.2% 0.891 0.867 0.859 19.0% 83.0%
Weighted Token Decoding 18.9% 0.865 0.846 0.844 17.3% 71.0% 29.6% 0.855 0.841 0.834 14.4% 72.3%
Mirostat Decoding 20.0% 0.918 0.903 0.901 12.2% 70.0% 31.2% 0.901 0.882 0.875 11.7% 79.1%
Rejection Sampling 19.8% 0.904 0.885 0.883 16.7% 72.0% 23.8% 0.890 0.865 0.857 19.5% 82.7%
Self-Debias Decoding 22.9% 0.895 0.879 0.878 14.3% 68.9% 26.0% 0.879 0.854 0.847 17.6% 83.9%
Local-Explanation Guard 20.2% 0.865 0.844 0.842 17.9% 73.5% 26.5% 0.870 0.847 0.836 20.1% 81.8%
Llama-3-8B-Instruct
Baseline 14.5% 0.855 0.837 0.836 7.0% 21.9% 0.847 0.834 0.828 4.0%
Self-Awareness Prompt 15.2% 0.909 0.881 0.878 22.9% 83.2% 21.0% 0.889 0.864 0.855 19.7% 85.5%
Chain-of-Thought 15.3% 0.909 0.881 0.878 22.6% 82.7% 22.5% 0.890 0.865 0.856 19.2% 85.6%
Cloze-Style Prompt 15.2% 0.909 0.881 0.878 21.9% 82.4% 21.9% 0.891 0.866 0.856 20.2% 85.3%
Cognitive Counterfactual Sim. 14.7% 0.909 0.881 0.878 22.5% 82.7% 21.8% 0.890 0.866 0.856 19.0% 84.8%
Self-Help Debias 15.2% 0.909 0.881 0.879 21.8% 83.0% 21.7% 0.891 0.865 0.856 19.2% 85.4%
Partial Summaries Ensemble 16.5% 0.890 0.871 0.870 15.4% 72.7% 28.5% 0.874 0.858 0.849 14.1% 79.3%
Weighted Summaries 15.1% 0.925 0.904 0.902 15.1% 77.0% 26.6% 0.880 0.866 0.855 9.0% 81.8%
Attention-Sort Re-ordering 16.9% 0.879 0.858 0.857 17.5% 74.7% 25.0% 0.869 0.845 0.831 20.7% 85.0%
Position-Invariant Shuffle 21.2% 0.899 0.888 0.887 11.5% 64.4% 27.0% 0.861 0.854 0.847 8.0% 66.3%
Forced Balanced Coverage 15.6% 0.900 0.878 0.877 19.5% 74.2% 23.8% 0.862 0.839 0.832 18.1% 81.3%
Weighted Token Decoding 13.6% 0.871 0.851 0.852 18.4% 70.0% 24.9% 0.852 0.831 0.823 18.1% 77.3%
Mirostat Decoding 13.8% 0.916 0.896 0.895 15.5% 75.8% 31.3% 0.900 0.875 0.867 18.3% 83.0%
Rejection Sampling 14.3% 0.911 0.890 0.889 18.9% 73.9% 26.8% 0.872 0.849 0.842 18.5% 80.0%
Self-Debias Decoding 16.0% 0.905 0.883 0.882 18.8% 75.1% 25.5% 0.879 0.855 0.848 18.2% 82.8%
Local-Explanation Guard 16.1% 0.859 0.837 0.836 18.9% 74.6% 25.9% 0.852 0.830 0.822 18.3% 79.1%
Phi-3-mini-4k-Instruct
Baseline 34.5% 0.836 0.822 0.823 5.4% 26.2% 0.828 0.827 0.823 4.9%
Self-Awareness Prompt 25.5% 0.882 0.866 0.867 14.5% 69.5% 24.9% 0.850 0.832 0.826 16.6% 73.7%
Chain-of-Thought 23.9% 0.884 0.868 0.868 14.1% 69.7% 26.0% 0.851 0.832 0.827 15.7% 73.5%
Cloze-Style Prompt 24.7% 0.883 0.868 0.868 13.6% 67.7% 25.9% 0.850 0.832 0.826 16.4% 73.5%
Cognitive Counterfactual Sim. 22.8% 0.877 0.862 0.862 14.4% 67.7% 25.5% 0.850 0.832 0.826 16.2% 73.5%
Self-Help Debias 22.8% 0.879 0.863 0.864 14.3% 68.5% 25.1% 0.851 0.832 0.826 16.0% 75.5%
Partial Summaries Ensemble 28.0% 0.893 0.880 0.875 11.5% 71.3% 27.7% 0.858 0.845 0.836 11.4% 73.9%
Weighted Summaries 22.6% 0.920 0.904 0.899 11.2% 78.1% 27.8% 0.880 0.871 0.863 7.4% 72.1%
Attention-Sort Re-ordering 33.1% 0.887 0.877 0.876 9.9% 66.8% 28.1% 0.837 0.823 0.818 14.5% 67.9%
Position-Invariant Shuffle 28.2% 0.877 0.866 0.865 9.3% 67.6% 28.2% 0.837 0.834 0.825 7.7% 63.0%
Forced Balanced Coverage 31.5% 0.901 0.891 0.893 7.6% 61.5% 29.9% 0.844 0.831 0.830 11.8% 66.3%
Weighted Token Decoding 25.0% 0.860 0.849 0.854 9.8% 56.9% 28.9% 0.841 0.829 0.828 14.4% 63.0%
Mirostat Decoding 33.2% 0.891 0.883 0.884 4.2% 58.8% 34.3% 0.854 0.845 0.843 8.7% 61.3%
Rejection Sampling 22.7% 0.885 0.876 0.878 8.2% 60.5% 28.3% 0.853 0.836 0.835 14.6% 66.8%
Self-Debias Decoding 28.8% 0.894 0.885 0.888 7.1% 56.4% 31.3% 0.844 0.832 0.834 11.3% 60.0%
Local-Explanation Guard 22.5% 0.886 0.875 0.879 9.4% 59.7% 27.3% 0.841 0.828 0.826 12.4% 65.8%
Table 9: Coverage, framing-change, and primacy-bias metrics for Amazon Reviews and MediaSum News. Metrics: $\varphi_{\mathrm{frame}}$ (framing-change fraction, $\downarrow$), $\bar{s}_{b},\bar{s}_{m},\bar{s}_{e}$ (cosine similarity with the first/middle/last third of the source, $\uparrow$), $\psi_{\mathrm{pri}}$ (share of summaries whose similarity to the beginning exceeds that to the middle by $>5\%$, $\downarrow$), and $\rho_{\text{pri}}$ (share of summaries exhibiting primacy bias, $\downarrow$).
Appendix E Prompt Templates Used in Our Experiments
Table 10 lists the exact prompt instruction templates supplied to each model under all the experimental settings.
Prompt Strategy Prompt
Summarization and General Bias Mitigation Prompts
Self-Awareness Prompting “You are an unbiased summarizer. Be mindful not to introduce any framing bias or omit the middle. Preserve the original sentiment.
Please summarize the following text: [DOCUMENT_TEXT]
FINAL_SUMMARY:”
Chain-of-Thought (Summarization) “Please read the text below carefully. Then break down the text into beginning, middle, and end, describing each portion in detail. After that, produce a final summary. Use the following format:
BEGIN_ANALYSIS: [describe the beginning]
MIDDLE_ANALYSIS: [describe the middle]
END_ANALYSIS: [describe the end]
FINAL_SUMMARY: [your final concise summary]
Text: [DOCUMENT_TEXT]
Cloze-Style Prompt “Fill the blanks for each part of the text:
BEGIN: ____
MIDDLE: ____
END: ____
Text: [DOCUMENT_TEXT]
When you fill them in, finally produce:
FINAL_SUMMARY:”
Cognitive Counterfactual Simulation (Final Revision) “Original text: [DOCUMENT_TEXT]
Initial Draft: [DRAFT_SUMMARY]
Here are the potential biased deviations: [LIST_OF_SIMULATED_BIAS_DEVIATIONS]
Please produce a final summary that avoids these biases.
FINAL_SUMMARY:”
Self-Help Debias (Rewrite Stage) “Rewrite the following summary to remove any bias, preserve the original sentiment, and ensure equal coverage of beginning, middle, and end: [DRAFT_SUMMARY]
FINAL_SUMMARY:”
Partial Summaries Ensemble (Merge Prompt) “Combine the following partial summaries into one coherent summary:
[PARTIAL_SUMMARY_1]
[PARTIAL_SUMMARY_2]
FINAL_SUMMARY:”
Weighted Summaries (Chunk Prompt) “Summarize this portion in about [PORTION_TOKEN_BUDGET] tokens:
[CHUNK_TEXT]
FINAL_SUMMARY:”
Attention-Sort Re-ordering (Final Prompt) “CONTEXT:
Segment 1: [SORTED_SEGMENT_1_TEXT]
Segment 2: [SORTED_SEGMENT_2_TEXT]
TASK: Summarize:
FINAL_SUMMARY:”
Position-Invariant Shuffle (Prompt) “The text is out of order; please summarize it fully: [SHUFFLED_DOCUMENT_TEXT]
FINAL_SUMMARY:”
Hallucination Mitigation Prompts
Baseline “You are a fact-checking AI. For each statement, predict if the event has occured or not and respond with ONLY ’True’ if it’s factual or ’False’ if it’s not. Do not provide any explanation. Respond with TRUE or FALSE only and nothing else.”
Prompt Calibration (CoT) “You are a fact-checking AI. Follow these strict guidelines:
- Respond with ’True’ ONLY if:
* You are 100% certain the statement describes an event that has occurred
* You have clear knowledge of the event from your training data
* You can identify specific details confirming the event happened
- Respond with ’False’ if:
* You have ANY doubt about the statement’s accuracy
* The statement contains ANY misleading or incorrect information
* You lack reliable information about the statement
Respond with EXACTLY one word: ’True’ or ’False’. No explanations.”
Knowledge Boundary Awareness “You are a fact-checking AI with knowledge up to {knowledge_cutoff}. For each statement, predict if the event has occurred or not. We check your decision-making abilities on futuristic events based on current data. Respond with only ’True’ or ’False’ for each news event.
If the event is scheduled to occur after {knowledge_cutoff}, make a good prediction based on your knowledge up to {knowledge_cutoff}.
RESPOND WITH EXACTLY ONE WORD: ’True’ or ’False’.”
Epistemic Tagging “You are a fact-checking AI. Analyze this statement and respond with:
- ’True [High Confidence]’ or ’True [Low Confidence]’
- ’False [High Confidence]’ or ’False [Low Confidence]’
Use these epistemic tags to guide your reasoning:
[Certain] = Information you know with high confidence based on your training data
[Uncertain] = Information you’re less confident about or may be outside your knowledge
[Reasoning] = Step by step analysis to determine veracity
[Conclusion] = Your final determination with confidence level
Respond ONLY with one of the four exactly formatted options listed above and nothing else.”
Table 10: Prompt strategies and the prompts used in our experiments.