### v2

# Quantifying Cognitive Bias Induction in LLM-Generated Content

Abstract
Large language models (LLMs) are integrated into applications like shopping reviews, summarization, or medical diagnosis support, where their use affects human decisions. We investigate the extent to which LLMs expose users to biased content and demonstrate its effect on human decision-making. We assess five LLM families in summarization and news fact-checking tasks, evaluating the consistency of LLMs with their context and their tendency to hallucinate on a new self-updating dataset. Our findings show that LLMs expose users to content that changes the context’s sentiment in 26.42% of cases (framing bias), hallucinate on 60.33% of post-knowledge-cutoff questions, and highlight context from earlier parts of the prompt (primacy bias) in 10.12% of cases, averaged across all tested models. We further find that humans are 32% more likely to purchase the same product after reading a summary of the review generated by an LLM rather than the original review. To address these issues, we evaluate 18 mitigation methods across three LLM families and find the effectiveness of targeted interventions.
Quantifying Cognitive Bias Induction in LLM-Generated Content
Abeer Alessa, Param Somane*, Akshaya Lakshminarasimhan* Julian Skirzynski, Julian McAuley, Jessica Echterhoff UC San Diego {aalessa, psomane, athenkarailakshminar, jskirzynski, jmcauley, jechterh}@ucsd.edu
11footnotetext: Equal contribution.
1 Introduction
LLMs perform well across numerous tasks Albrecht et al. (2022), such as content summarization Laban et al. (2023), translation Elshin et al. (2024), sentiment analysis Zhang et al. (2024). In many of these tasks, humans rely on LLMs for daily decision-making support Rastogi et al. (2023); Li et al. (2022) in expert contexts such as writing policy documents Choi et al. (2024) or summarizing medical documents (Spotnitz et al., 2024). However, models have been shown to inherit or exhibit several societal biases Zhao et al. (2018); Nadeem et al. (2020); Liang et al. (2021); He et al. (2021), e.g., favoring specific genders Zhao et al. (2018) and ethnicities Caliskan et al. (2017), or learn similar decision-making-patterns as cognitively biased humans Echterhoff et al. (2024).
Any bias within the model can affect how textual context is processed for a human task and introduce biased content to users. Our work specifically focuses on qualifying and quantifying the extent to which biased content is induced by LLMs to users, as visualized in Figure 1 with examples shown in Figure 2. We summarize our contributions as follows:
Figure 1: LLMs alter information from a source text when performing a task for the user (e.g., when an LLM summary has a different sentiment compared to the original text). This model behavior introduces biased content to humans and can hence affect their decision-making. We evaluate how LLMs highlight source content, leading to exposure/primacy bias for users, how LLMs reframe sentiments, leading to framing bias for humans, and how LLMs hallucinate, leading to authority/confirmation bias.
Figure 2: When LLMs process context for users to consume, they may change its content, e.g., by changing its sentiment or omitting some relevant parts. Exposure to altered content may elicit cognitive biases, such as positive framing bias or primacy bias, and subsequently lead humans to make different decisions than they would if they saw the original text.
  1. 1.
    We analyze the extent of content alteration that LLMs introduce in summarization and news fact-checking tasks. We find that LLMs alter the sentiment or framing of a text in 26.42% of summaries, disproportionately focus on the beginning of a source text in 10.12% of summaries, and hallucinate in 60.33% of post-knowledge-cutoff news summaries for real-world fact-checking questions, averaged across all tested models.
  2. 2.
    We find that human users are affected by biases introduced by LLMs. Participants in our human-subject study were 32% more likely to purchase a product when they read its positively-framed LLM summary than when they read the original review.
  3. 3.
    We provide a dataset that enables analysis of model hallucination after a model’s cut-off date, which is self-updating to remain relevant for future models and prevent leakage into training data. At the time of writing this paper, it has 8,554 data points [^1] https://huggingface.co/datasets/sparklessszzz/NewsLensSync .
  4. 4.
    We evaluate 18 distinct mitigation methods and highlight which ones were effective for specific types of content alterations.
2 Related Work
2.1 Social and Representation Biases in LLMs
Previous work has demonstrated the presence of various biases in LLMs. Primacy and recency bias in LLMs have been shown in question-answering settings, where models tend to be more accurate when the answer is presented earlier or later in the documents Liu et al. (2024); Peysakhovich and Lerer (2023); Wang et al. (2023); Zhang et al. (2023); Eicher and Irgolič (2024). When presented with options, LLMs have been shown to exhibit sensitivity to their order Zheng et al. (2023a), and framing Leto et al. (2024). In framing bias, alterations in the tone, e.g., from positive to negative, can change the model’s decisions Jones and Steinhardt (2022); Echterhoff et al. (2024). In contrast to prior work, we quantitatively assess the extent to which LLM responses are biased.
2.2 Cognitive Effects of LLM-Generated Content on Humans
When given a specific task, a model can alter, omit, or add information to produce the final output, which can affect users in several ways Wester et al. (2024); Choi et al. (2024). Previous work demonstrates that LLMs can influence users by reinforcing their existing opinions Sharma et al. (2024), increasing confirmation bias, or influencing political decision-making Fisher et al. (2024). The style in which LLMs present advice can significantly influence users’ perceptions of its usefulnessWester et al. (2024). Altered headlines can introduce a framing effect on individuals in collective decision-making Abels et al. (2024). Choi et al. (2024) found that even when experts use LLMs, they tend to adopt the suggestions with fewer revisions.
Our work quantifies the extent of bias the model generates in its output compared to its context by LLMs, and finds that these content changes influence human purchasing decisions in a human-subject study.
3 Background
Humans are prone to mental shortcuts that negatively influence rational decision-making, known as cognitive biases. These mental shortcuts are often amplified by simple changes of text, framing, or highlighting that could be introduced by an LLM. We focus on three cognitive biases in our evaluation setup, which we briefly introduce here.
Framing bias refers to changes in human decisions based on how a problem is presented. Individuals make different choices based on the wording of questions, even when the underlying options remain unchanged Tversky and Kahneman (1981). Based on prior work, we assume that a change in framing affects users and should be avoided because it can influence their decisions, e.g., when comparing the framing of a factual context with its summarization.
Primacy Bias. Primacy bias occurs when humans prioritize information encountered first. Murphy et al. (2006). This bias skews decision-making based solely on the order in which information is encountered Glenberg et al. (1980). When LLMs exhibit primacy bias Echterhoff et al. (2024), they may filter or highlight information, leading to exposure bias. We assume that systematically highlighting any part of the context, regardless of its position, can influence users and should be avoided.
Confirmation and Authority Bias. Individuals favor information confirming their existing beliefs or expectations Nickerson (1998). We assume that if a user prompts a model with belief-consistent input and the model subsequently hallucinates, this can reinforce the user’s prior beliefs, leading to biased or misinformed decisions. Authority bias causes individuals to trust information from perceived experts or authoritative sources Milgram (1963); Cialdini (2007); Gültekin (2024). Research on human-AI interaction suggests that similar dynamics apply when interacting with artificial systems perceived as experts, which can create challenges when verifying facts Glickman and Sharot (2025). For example, if a user asks an LLM to verify whether a particular legislation was passed, a user with strong prior beliefs may exhibit confirmation bias, while others may exhibit authority bias.
4 Quantifying Content Alteration in LLMs Responses
To capture unintended content alterations introduced by LLMs, we propose the following metrics, motivated by the aforementioned human cognitive biases. We focus on content alteration introduced through changes in content framing, content filtering/highlighting, and content truthfulness.
Framing Effects in Model-Generated Content
We test whether a model changes the sentiment of its summary compared to the source context. For example, a model may shift the summary sentiment to positive, whereas the source context was neutral. To evaluate these changes, we classify the framing of the source context $f_{c}$ (positive, negative, or neutral) before summarization, as well as the framing of the generated summary $f_{s}$. [^2] We use GPT-o4-mini to classify the framing of the text, after evaluating different models for this task. Details are provided in Appendix A.1.
We consider a model consistent if $f_{c}$ equals $f_{s}$. We define the bias-inducing score as the framing-change fraction, denoted by $\varphi_{\mathrm{frame}}$ as follows:
$\varphi_{\mathrm{frame}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl(f_{c}^{(i)}\neq f_{s}^{(i)}\bigr),$ (1)
where $\mathbb{I}(\cdot)$ is the indicator function that returns one if the condition is true and zero otherwise.
Primacy Effects in Model-Generated Content
We assess whether a model emphasizes a specific part of its context. For example, a model might focus solely on the beginning of an Amazon user review that starts with praise but concludes with criticism. We divide the source context into three equal segments and calculate the cosine similarity, using embeddings Wang et al. (2024), between each segment and the summary. We denote the similarities of the summary and the beginning as $s_{b_{i}}$, the middle as $s_{m_{i}}$, and the end as $s_{e_{i}}$ for each instance $i$. We define biased examples as those where the similarity to the beginning exceeds the similarity to the middle by at least a threshold $\alpha$. This comparison is designed to isolate the primacy effect (over-reliance on initial content) from potential recency effects (over-reliance on final content), which can be measured separately. We define this primacy bias score as:
$\psi_{\mathrm{pri}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left(s_{b_{i}}>s_{m_{i}}+\alpha\right)$ (2)
with $\alpha=0.05$ for this study. A higher score indicates that a greater proportion of summaries focus more on the beginning of the source content than on the middle.
Hallucination in Model-Generated Content
We quantify whether a model generates hallucinated content when asked to fact-check news content past its knowledge cutoff. For example, a model may incorrectly state that an event never occurred, even if it did occur after the model’s training was completed. We prompt a model to verify the truth of specific news items. We test the accuracy of true, real-world news (Actual News Accuracy) and its falsified (negated) counterparts (Falsified News Accuracy). We evaluate the strict accuracy as correctly identifying both true and falsified news, with the paired instances of true news and its falsified counterparts:
$s_{H}^{T}(D_{j})=\frac{1}{|P_{j}|}\sum_{(x_{i}^{t},x_{i}^{f})\in P_{j}}\mathbb{I}(\hat{y}_{i}^{t}\land\neg\hat{y}_{i}^{f})$ (3)
where $P_{j}$ is the set of paired instances in dataset $D_{j}$, $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{f}$ are the model’s predictions for the true and falsified versions of the same news item, respectively. $T$ represents the time horizon (pre- or post-model-knowledge-cutoff).
We analyze the disparity between pre-cutoff and post-cutoff performance:
$\delta_{H}=|s_{H}^{\mathit{pre}}-s_{H}^{\mathit{post}}|$ (4)
A larger $\delta_{H}$ indicates a higher fraction of generated incorrect/hallucinated information.
Models
We evaluate six open- and closed-source language models from five families to ensure our findings are generalizable across architectures. Our selection includes small open-source models (Phi-3-mini-4k-Instruct, Llama-3.2-3B-Instruct, Qwen3-4B-Instruct), a medium-sized model (Llama-3-8B-Instruct), a large open-source model (Gemma-3-27B-IT), and a closed-source model (GPT-3.5-turbo). Details on all models, model selection per task and our hyperparameter settings are provided in Appendix A.2. Appendix Table A2 shows knowledge cutoff dates for each model.
5 Datasets
News Interviews
News summaries should cover all key information, regardless of where it appears, and remain neutral, matching the tone of the original content. We use MediaSum, a dialogue summarization dataset comprising media interviews from NPR and CNN Zhu et al. (2021), to measure framing and primacy effects. We randomly select 1,000 data points with $\leq$ 4,000 tokens each, ensuring that the full content fits within a model’s context window. The descriptive statistics of text lengths are provided in the Appendix in Table A3.
Product Reviews
Previous work suggests that changing the framing of the product influences consumer purchase decisions Wei et al. (2024). To explore this effect in the context of LLM outputs, we use the Amazon Reviews dataset  Ni et al. (2019). This dataset comprises customer reviews from Amazon across various product categories, providing insights into product quality, usability, and customer satisfaction Ni et al. (2019). We sample 1,000 random examples from the Electronics category, ensuring token lengths of $\leq$ 4,000 to maintain full-context input for the model. Prior work notes that primacy bias is more observed in longer context windows Liu et al. (2024). Because most Amazon reviews are short, we evaluate the 1,000 longest reviews (by token length). Appendix Table A3 reports the descriptive statistics of the resulting dataset.
News Hallucination
We provide a novel dataset to quantify how models discern factual information across temporal boundaries and in both true and negated versions of news articles.
Pre-Knowledge Cutoff Data: To evaluate models on content before their knowledge cutoff, we sample 2,700 random data points from the “All the News” dataset all (2020), which includes articles from 27 major U.S. publications between 2016 and 2020.
Post-Knowledge Cutoff Data: To assess model performance on unseen events, we introduce a self-updating dataset, NewsLensSync, designed to evaluate hallucinations beyond model knowledge cutoffs. It is built using NewsAPI new , accessing live and historical content from 150,000 publishers globally. We restrict our queries to the keyword "politics," English-only articles, and dates after April 2025, being past all model knowledge cutoffs in our study.
Counterfactual versions are generated using a transformer model to falsify the descriptions of the new articles Anschütz et al. (2023). This results in paired authentic/falsified versions of the same news items which are manually validated for semantic consistency, producing matched authentic/negated pairs. Each data point includes a publication timestamp for filtering according to model cutoff dates. Further information on our self-updating dataset is described in Appendix A.3
6 Mitigation Strategies
We introduce targeted interventions designed to ensure balanced content coverage, reduce misleading emphasis, and maintain accuracy in model outputs. We analyze a total of 18 mitigation methods, but describe only the most promising in this section. All other mitigation methods are shown in Appendix A.6.
6.1 General Purpose Mitigation
Self-Awareness Prompt (SA).
Simple awareness prompts can mitigate multiple biases Mair et al. (2014); Echterhoff et al. (2024). To use this prompt-based mitigation in our summarization setting, we use the following prompt:
“You are an unbiased summarizer; be mindful not to omit coverage of any portion of the text or alter its sentiment. Do not incorporate framing bias.”
Chain-of-Thought Prompting (CoT).
Chain-of-Thought prompting encourages the model to decompose complex problems by generating a series of intermediate reasoning steps before producing a final answer Wei et al. (2023). In the context of summarization, we guide the model to articulate steps for processing different parts of the input text. For the question-answering setting, we explicitly define when to return specific responses during fact-checking.
6.2 Primacy Effect Mitigation
We group interventions into four categories: prompt-level cues, chunk-based summarization, attention/order re-ranking, and decoding-time control. Each category targets the tendency of LLMs to overemphasize early context (primacy) or alter the source sentiment. All methods we experimented with are further described in Appendix A.6.
Weighted Summaries is a chunk-based approach that ensures balanced representation of the source text. The algorithm first splits the document into three segments (beginning, middle, and end). It then generates a partial summary for each segment independently, constraining the output length for each part according to a pre-allocated token budget (33%, 34%, and 33% of the total desired summary length). These partial summaries are then combined to form the final output.
In Mirostat Decoding, the model’s temperature is dynamically adjusted during generation via a feedback loop (formula in Appendix Table A5). This ensures the output text maintains a consistent level of unpredictability, which prevents the model from focusing too narrowly on initial sentences and encourages broader coverage of the source text. We use a target entropy of $\mu^{\star}\!=\!2.0$ in our runs, similar to Basu et al. (2021).
6.3 Framing Effect Mitigation
Weighted Token Decoding directly influences the model’s token selection during the generation process. At each step, we modify the output probabilities by adding a logarithmic weight to the logits of specific tokens before sampling Liu et al. (2021); Dathathri et al. (2020). In particular, we down-weight tokens associated with negative sentiment to reduce framing shifts, as detailed in Appendix A.6.
6.4 Hallucination Mitigation
We analyse Knowledge Boundary Awareness, which explicitly defines the model’s knowledge cutoff date in the context to establish clear temporal limits on what the model can reasonably know. Epistemic Tagging requires models to express confidence levels alongside factual assertions, forcing self-evaluation of fact-checking abilities (i.e., the model additionally responds with a high or low confidence level for each response) Lin et al. (2022).
Hallucination mitigation strategies are detailed in Appendix A.6 and prompt details provided in Appendix A.8.
Amazon Reviews MediaSum News
Method $\varphi_{\mathrm{frame}}$ $\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\varphi_{\mathrm{frame}}$ $\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$
GPT-3.5-turbo 16.0% 0.848 0.826 0.825 7.6% 24.8% 0.840 0.823 0.820 6.1%
Llama-3.2-3B-Instruct 14.9% 0.860 0.842 0.840 7.4% 22.1% 0.851 0.837 0.832 5.1%
Llama-3-8B-Instruct 14.5% 0.855 0.837 0.836 7.0% 21.9% 0.847 0.834 0.828 4.0%
Phi-3-mini-4k-Instruct 34.5% 0.836 0.822 0.823 5.4% 26.2% 0.828 0.827 0.823 4.9%
Qwen3-4B-Instruct 27.8% 0.921 0.896 0.893 19.6% 45.9% 0.898 0.879 0.869 13.0%
Gemma-3-27B-IT 26.5% 0.924 0.896 0.892 23.7% 41.9% 0.902 0.879 0.868 17.6%
Table 1: Models introduce framing changes ($\varphi_{\mathrm{frame}}$) and filter content based on position ($\psi_{\mathrm{pri}}$). $\varphi_{\mathrm{frame}}$ is the proportion of examples where the framing changed from the original to the summary (lower is better). $\bar{s}_{b}$, $\bar{s}_{m}$, and $\bar{s}_{e}$ represent the average cosine similarity between the summary and the beginning, middle, and end thirds of the source text, respectively (higher is better). $\psi_{\mathrm{pri}}$ is the percentage of examples with primacy bias as defined in Equation 2 (lower is better). All differences between $\bar{s}_{b}$ and $\bar{s}_{m}$ are statistically significant ($p<.001$).
7 Results
7.1 LLMs Alter Framing During Summarization
[…]
MediaSum GPT-3.5-turbo 2.9% 8.6% 0.1% 0.1% 1.9% 11.2%
Llama-3-8B-Instruct 2.6% 8.0% 0.0% 0.1% 2.3% 8.9%
Llama-3.2-3B-Instruct 2.0% 7.7% 0.1% 0.1% 2.3% 9.9%
Phi-3-mini-4k-Instruct 2.6% 7.3% 0.1% 0.7% 2.4% 13.1%
Qwen3-4B-Instruct 7.6% 31.4% 0.7% 0.3% 2.0% 2.9%
Gemma-3-27B-IT 5.2% 30.8% 0.8% 0.1% 1.6% 3.0%
Amazon GPT-3.5-turbo 2.6% 0.5% 0.2% 1.4% 7.4% 3.9%
Llama-3-8B-Instruct 2.3% 0.6% 0.7% 1.4% 6.7% 2.8%
Llama-3.2-3B-Instruct 1.8% 0.6% 0.5% 2.1% 6.2% 3.7%
Phi-3-mini-4k-Instruct 1.6% 0.8% 1.8% 2.2% 21.2% 6.9%
Qwen3-4B-Instruct 15.3% 3.9% 1.5% 1.4% 1.5% 1.3%
Gemma-3-27B-IT 14.4% 4.2% 1.8% 1.3% 1.2% 0.7%
[…]
We observe that models alter the framing of summaries, compared to the source content, in both media interviews and the Amazon dataset (Table 1). GPT-3.5-turbo and Llama models exhibit a framing change fraction between 14.5% and 16%, whereas Gemma-3-2-27B-IT, Qwen3-4B-Instruct, and Phi-3-mini-4k-Instruct show framing changes between 26.5% and 45.9%. In Table 3, we show an example of a framing change between a summary and the ground-truth context. Table 2 provides a detailed evaluation of framing alterations (e.g., from neutral to positive, or positive to negative). In the MediaSum news interviews, most framing shifts occur between neutral and negative, ranging from 7.3% to 8.6% for GPT-3.5-turbo and Llama models, and reaching 30.8% to 31.4% for Qwen3-4B-Instruct and Gemma-3-27B-IT. Shifts from negative to neutral range from 2.9% to 13.1% across models. In Amazon reviews, the most common shift is from positive to neutral, ranging from 6.2% to 7.4% for the Llama and GPT-3.5-turbo models, and 21.2% for Phi-3-mini-4k-Instruct. Qwen3-4B-Instruct and Gemma-3-27B-IT show higher rates of neutral-to-positive shifts (14.4% to 15.3%) compared to other models, suggesting different framing patterns across model families.
Original Review Summary
“This is a fun, useful tablet for the price. I love that you can make all different homepages. Love that you can pick wallpaper which you can’t do with Kindle Fires, which are the only other tablets I’ve owned. Of course, if you want a tablet mainly for books, Kindle is the tablet you want. I also love all the apps you can get with the Samsung. What I don’t like is the 8 GB storage. I was under the impression that if you got the 32 GB extra, it would integrate with the tablet rather than just be like a drive for MP3s and such. 8 gigs is barely enough for the apps you need. Also, the manual tells you nothing about how to add things to the extra drive. Had to go to YouTube. My tablet has connection issues from time to time. I still really enjoy it most of the time. It can be super fast and plays things well from the cloud. Very light and attractive. I am returning the tablet. 8 GB is just not enough. Extra storage is cumbersome to use and doesn’t work half of the time. Not easy like the cloud. Speakers unimpressive. I am disappointed because I thought I would like the Samsung. Also a battery hog. Will try the Nexus 7.” “I enjoyed the Samsung tablet for its customization options and app variety, but the 8 GB storage was insufficient. The extra storage was cumbersome and unreliable. Connection issues and battery drain were also drawbacks. I will be returning it and trying the Nexus 7 instead.”
Table 3: Example of framing shift: from neutral (original text) to negative (summary by GPT-3.5-turbo).
7.2 LLM Summaries have Imbalanced Context Coverage
We find that, on average, summaries align more closely with the beginning of the source text than with the middle and end sections (Table 1). For example, across all models and datasets, the average similarity to the beginning segment is consistently higher than the similarity to the middle and end segments. Paired t-tests between the beginning-summary similarity ($s_{b}$) and middle-summary similarity ($s_{m}$) are statistically significant across all evaluated models and datasets ($p<.001$). The detailed test statistics are provided in Appendix A.5. We further quantify this coverage imbalance using our primacy bias score. For the Amazon Reviews dataset, $\psi_{\mathrm{pri}}$ ranges from 5.4% to 23.7%. Similarly, MediaSum summaries exhibit $\psi_{\mathrm{pri}}$ scores between 4.0% and 17.6%, which shows primacy bias in structured news content.
7.3 LLMs Hallucinate on Post-Knowledge-Cutoff Data
Table 4 shows a consistent decrease in strict accuracy $(s_{H}^{T})$ from pre-cutoff to post-cutoff data across all models. Llama-3-8B-Instruct declined from 26% to 21% ($\delta_{H}=5\%$), Llama-3.2-3B-Instruct dropped from 19% to 13% ($\delta_{H}=6\%$), and Phi-3-mini-4k-Instruct showed a $\delta_{H}=4\%$ decline.
The consistently low strict accuracy, compared to actual news and falsified news accuracy, highlights a critical limitation: the persistent inability to reliably differentiate fact from fabrication. Strong performance on isolated real or falsified news does not guarantee that models can reliably distinguish truth from falsehood.
Models Prompt Strategy Pre-cutoff data Post-cutoff data
Actual News Acc. Falsified News Acc. Strict Acc. $(s_{H}^{T})$ Actual News Acc. Falsified News Acc. Strict Acc. $(s_{H}^{T})$
Llama-3-8B-Instruct Baseline 74% 46% 26% 75% 42% 21%
Prompt Calibration (CoT) 39% 85% 27% 36% 84% 22%
Knowledge Boundary Aware 65% 54% 25% 50% 63% 18%
Epistemic Tagging 80% 35% 18% 78% 37% 19%
Llama-3.2-3B-Instruct Baseline 48% 58% 19% 35% 72% 15%
Prompt Calibration (CoT) 48% 56% 17% 46% 63% 15%
Knowledge Boundary Aware 25% 87% 15% 15% 94% 10%
Epistemic Tagging 51% 68% 26% 53% 64% 23%
Phi-3-mini-4k-Instruct Baseline 14% 97% 12% 9% 98% 8%
Prompt Calibration (CoT) 3% 99% 3% 3% 99% 3%
Knowledge Boundary Aware 3% 99% 3% 2% 99% 2%
Epistemic Tagging 45% 76% 31% 53% 70% 29%
Table 4: Accuracies on real news items and their falsified counterparts, evaluated before and after each model’s training cutoff date. Strict Accuracy $s_{H}^{T}(D_{j})$ counts predictions where the model gets both members of each (True, False) pair correct.
7.4 Special Biases Require Special Mitigations
We discuss trade-offs among mitigation methods for different biases, and how their effectiveness varies depending on the model and corpus. For example, techniques aimed at reducing position-based salience (e.g., primacy bias) often work by redistributing attention across the input. However, this can disrupt the model’s ability to maintain consistent sentiment in its responses. All definitions of the 18 mitigation approaches we tested are in Appendix A.6.
Method $\bar{s}$ $\psi_{\mathrm{pri}}$% $\varphi_{\mathrm{frame}}$% §
Baseline 0.843 7.0% 14.5%
Weighted Summaries 0.910 15.1% 15.1%
Mirostat Decoding 0.902 15.5% 13.8%
Pos.-Invariant Shuffle 0.891 11.5% 21.2%
Weighted Token Decoding 0.858 18.4% 13.6%
Self-Awareness Prompt 0.889 22.9% 15.2%
Table 5: Key mitigation results on Amazon Reviews for Llama-3-8B-Instruct. Mean content-coverage similarity $(\bar{s}_{b}+\bar{s}_{m}+\bar{s}_{e})/3$. Primacy-bias score. §Framing-change rate. Lower is better for $\psi_{\mathrm{pri}}$ and $\varphi_{\mathrm{frame}}$. Behavior on other corpora/models follows a similar qualitative pattern; see full results in Table A6.
7.4.1 Primacy Effect and Coverage
The Weighted Summaries method, which assigns a fixed token budget to each text segment, consistently increases overall content coverage. For instance, with Llama-3-8B-Instruct on Amazon Reviews (Table 5), it increases average coverage similarity from a baseline of 0.843 to 0.910. However, this comes at the cost of an increased primacy bias score ($\psi_{\mathrm{pri}}$), which increases from 7.0% to 15.1%.
Mirostat Decoding also improves coverage (Table A6); it is the only method that reduces $\psi_{\mathrm{pri}}$ (primacy bias) on the smaller Phi-3 model by 1.2 percentage points (from 5.4% to 4.2%) but worsens it on larger models (e.g., from 7.0% to 15.5% on Llama-3-8B for Amazon).
Position-Invariant Shuffle is useful diagnostically because it helps confirm that the model relies on word order. By randomly shuffling sentences, we remove positional cues; the resulting change in output reveals the model’s sensitivity to input structure. While this method increases overall coverage (e.g., from 0.843 to 0.891 for Llama-3-8B on Amazon Reviews; Table A6), it degrades other metrics, increasing the primacy score $\psi_{\mathrm{pri}}$ to 11.5% and the framing change rate $\varphi_{\mathrm{frame}}$ to 21.2%, and comes at the risk of losing temporal information.
7.4.2 Framing Effect and Consistency
Weighted Token Decoding is the most effective method for mitigating framing bias. By downweighting negative lexemes during decoding, it achieves the lowest framing-change fraction ($\varphi_{\mathrm{frame}}$) for Llama-3-8B, reducing it by 0.9 percentage points to 13.6% (Table A6). However, this targeted intervention significantly worsens primacy bias, with $\psi_{\mathrm{pri}}$ increasing from 7.0% to 18.4%.
Prompt-only approaches: Prompts such as Self-Awareness or Chain-of-Thought, alter $\varphi_{\mathrm{frame}}$ by only a few percentage points in most settings. However, the Self-Awareness variant yields the best framing result on MediaSum for Llama-3-8B-Instruct, reducing $\varphi_{\mathrm{frame}}$ to 21.0% from a 21.9% (Table A6) baseline at a negligible computational cost.
7.4.3 Prompt Calibration and Epistemic Tagging improve Hallucination
Applying Prompt Calibration with Chain-of-Thought enhances strict accuracy for Llama-3-8B Instruct but shows limited utility for others. It improves the detection of falsified news by 40% for both pre- and post-cut-off data.
Knowledge Boundary Awareness improves Llama-3-8B-Instruct’s performance by 8% and 21% in falsified news detection in pre and post cut-off data, respectively, alongside maintaining its strict accuracy close to the baseline. For the other models, performance declines, despite observing a significant improvement in the detection of falsified news data.
Epistemic Tagging performs best across all models. It associates a confidence level during the fact-checking process (i.e., high or low) with its response about whether an event occurred (Table 6)
We observe an increase in strict accuracy for Phi-3-mini-4k-Instruct by 21% on post cut-off data and by 19% on pre cut-off data. The Llama-3.2-3B-Instruct also shows improved accuracy of 8% on post-cut-off data and 7% on pre-cut-off data, with balanced independent accuracies above 50% for real and falsified news data.
Models cutoff Actual News Falsified News
High Low High Low
Llama-3-8B-Instruct Pre 99.0% 1.0% 99.8% 0.2%
Post 98.3% 1.7% 99.9% 0.1%
Llama-3.2-3B-Instruct Pre 75.8% 24.2% 24.6% 75.4%
Post 66.7% 33.3% 83.3% 16.7%
Phi-3-mini-4k-Instruct Pre 93.8% 6.2% 86.2% 13.8%
Post 98.1% 1.9% 95.4% 4.6%
Table 6: Epistemic Tagging prediction Confidence levels (high, low) for different models across cutoff conditions.
8 Impact on Human Decision-Making
When LLMs modify the framing or emphasize certain aspects of the content they process, they can inadvertently shape human perception and decision-making when humans are exposed to the content. In this section, we empirically demonstrate that human decisions change when LLMs process the original text and introduce undesired changes to it.
Experiment Design
Participants choose between two manufacturers of a product (e.g., a headlamp) based on product reviews. We probe 1000 reviews for items in the home, electronics, and kitchen categories from the Amazon-Reviews-2023 dataset Hou et al. (2024) and generate summaries using GPT-3.5. To analyze the worst-case scenario of how LLM content alteration can influence decision-making, we use GPT-4 to identify product types with the most extreme framing changes and manually select products to show to participants. Specifically, we select 10 product pairs where one product review shifts from negative/neutral to positive after summarization, while the other maintains neutral framing.
Participants are divided into two conditions and view 10 product pairs, either as original reviews or summaries (5 of each format per participant, alternating between conditions). We measure manufacturer selection rates based on the framing of reviews. Additionally, participants are asked to provide their Willingness to Pay Mitchell and Carson (2013) (WTP) for the product from the manufacturer they picked. WTP describes the maximum price someone would pay, knowing the cost of another product. For example, if the participant selects manufacturer $M$ whose product is described by a positive summary, they learn that the product from manufacturer $W$ costs $\mathdollar X_{W}$ and decide how much more over $\mathdollar X_{W}$ they would pay to $M$. This measure enables us to quantify the strength of the change in behavior induced by altering the content by the LLM.
Hypothesis
We hypothesize that participants will be more likely to choose a manufacturer when they read a positively framed LLM summary of their product compared to when they read the original review (which is neutral or negative). Additionally, we posit that their Willingness to Pay will be higher when reading the positively framed summaries.
Participants
We recruit 72 participants on Prolific (32-40 per condition, English-speaking, 33 men, 39 women, aged 19-75). Participants are presented with instructions about the study and take a short comprehension quiz at the beginning of the study, before proceeding to the decision-making stage. Based on this quiz, we accept 70 submissions (3% rejection rate). On average, the experiment takes 32 minutes and pays \$4.83, which includes \$3 base pay and a performance bonus for selecting the option that matches the majority opinion (encouraging participants to make decisions they believe reflect typical consumer preferences).
Figure 3: Mean selection rates for manufacturers depending on whether participants read the original review of their product (neutral or negative framing) or its summary (positive framing). Error bars display 95% confidence intervals and stars denote significance levels (* p < 0.05, ** p < 0.01, *** p < 0.001).
Results
We see that humans make different decisions, depending on the framing of the text they are exposed to. Choice rates differ significantly between framing conditions ($\chi^{2}$ = 78, df = 1, $p<0.001$). Fig.˜3 shows that participants picked manufacturers when their products were described by summaries (positive framing) in 83.7% of the cases compared to 52.3% of the cases if these products were described by original reviews (neutral or negative framing). The effect is robust across products, but differs in strength (Fig.˜3). Additionally, we find that reading the better-framed summaries increases participants’ willingness to pay by 4.5% with respect to the original product price ($t=4.406,p<0.001$ according to the linear mixed-effects model). Together, these results demonstrate that exposing participants to summarized content significantly affects whether they would purchase a product and what price they would pay for it. This demonstrates that when LLM summarization alters the text’s framing, it can have measurable consequences on human behavior.
9 Conclusion
This paper quantifies content changes in summarization and news fact-checking tasks performed by LLMs. Averaged across all tested models, models introduce framing bias to the user in 26.42% of instances, and show primacy bias in 10.12% of the cases, where summaries disproportionately reflect content from the beginning of the original text. Users are exposed to hallucinations on post-knowledge-cutoff content in 60.33% of instances ( averaged across all tested models). Through our user study, we find that model bias (specifically, framing bias that occurs in review summarization) influences people’s decisions. We observe that a positive framing bias increases the likelihood that people will purchase a given product and enhances their perception of its value.
By evaluating 18 mitigation strategies, we find that each approach’s effectiveness depends on the model and the targeted bias. For example, Weighted Summaries and Mirostat Decoding show reduced framing changes and primacy biases, particularly in smaller and mid-sized models. Epistemic Tagging consistently improved factual reliability regarding post-knowledge-cutoff data for smaller models.
Our paper represents a step toward careful analysis and mitigation of content alteration introduced by LLMs to humans, and provides insight into its effects, aiming to reduce the risk of systemic bias in decision-making across media, education, and public policy.
10 Limitations
This study investigates the degree to which LLM outputs modify original content in ways that may impact human judgment. Most of our mitigation methods were designed with one specific issue in mind (e.g., framing, hallucinations, or primacy effects). In our future work, we intend to explore more general mitigation techniques that address overlapping biases simultaneously.
Our current primacy bias score uses a fixed, generic threshold to identify summaries that disproportionately focus on the beginning of the source text. While this threshold provides a consistent way to detect bias, our future work aims to explore ways to refine it. In addition, our evaluation relies on one LLM to assess another LLM’s outputs in the framing bias metric. Although commonly used, LLM-as-a-judge approaches have drawbacks. We use LLM-as-a-judge to the best of our knowledge, given current best practices.
Our work aims to encourage the NLP community to continuously test LLMs for biases that can influence users’ decision-making, which is why we introduce the self-updating NewsLensSync dataset. We urge researchers to exercise caution when framing and contextualizing their use of the data. We provide our data under the CC-BY-4.0 license.
Ethical Considerations.
Our analysis reveals that LLMs tend to focus more on content from the beginning of the context, exhibit hallucination of facts, and alter the text framing in their outputs. These alterations can affect the user and influence their decision-making. These findings highlight the risks associated with working with LLMs, even in everyday tasks.
Experiments. All experiments are run on NVIDIA RTX A6000 for open-source models and the official APIs for closed-source models with a fixed random seed. Our run time was approximately 360 GPU hours. All prompts details are included in Appendix A.8.
Appendix A Appendix
A.1 LLM as a framing judge
Previous work has shown that high-capability LLMs like GPT-4 can achieve over 80% agreement with human experts on question-answering tasks, with the level of agreement observed among humans themselves being 81% cite zheng2023judging. GPT-4 reached 85% agreement with human judgments, suggesting that LLMs can serve as effective proxies for human preferences Zheng et al. (2023b). Thus, we test three high-capability models to classify the framing in our study ( GPT-3.5-turbo, GPT-4-turbo, and GPT-o4-mini). On a sample of 500 randomly selected Amazon reviews, we use the user-provided ratings as ground truth to evaluate the models’ accuracy. We ask the models to rate the product based on the review, then we map ratings 1 and 2 as negative, 3 as neutral, and 4 and 5 as positive. The GPT-3.5-turbo, GPT-4-turbo, and GPT-o4-mini achieve accuracy of 0.89, 0.91, and 0.92, respectively. Based on this, we select GPT-o4-mini to serve as the framing judge. We prompt the model as follows:
“[TEXT]
Classify the framing of the text as Positive, Negative, or Neutral. Respond with the class label only.”
to obtain the text framing.
A.2 Model Setup and Hyperparameters
Our implementation uses the Hugging Face Transformers library via its official APIs. We employ Flash Attention 2 when available to enhance computational efficiency during inference. For consistency across experimental conditions, we maintain fixed generation parameters with temperature=0.01, do_sample=False, and max_new_tokens= 500 set to a constant with a fixed random seed. Table A1 presents the models we evaluate, their context windows, and the rationale for choosing them.
All aggregate statistics reported in the abstract are computed using a mean-of-means approach across all six models (GPT-3.5-turbo, Llama-3-8B-Instruct, Llama-3.2-3B-Instruct, Phi-3-mini-4k-Instruct, Qwen3-4B-Instruct, Gemma-3-27B-IT) and both datasets (Amazon Reviews and MediaSum News), where each model-dataset pair contributes one mean calculated from 1000 examples.
Model Context Window Reason
GPT-3.5-turbo 16k A closed-source, highly capable model that is cost-effective OpenAI (2024).
Llama-3-8B-Instruct 8k Outperforms competing models on both per-category win rate and average per-category score Grattafiori et al. (2024).
Llama-3.2-3B-Instruct 8k Lightweight yet powerful; offers multilingual text-generation abilities Meta AI (2024).
Phi-3-mini-4k-Instruct 4k Matches high-performance closed-source models on several key benchmarks Abdin et al. (2024).
Qwen3-4B-Instruct 256k A small, high-performing open-source model included to broaden architectural diversity Yang et al. (2025).
Gemma-3-27B-IT 128k A large, state-of-the-art open-source model included to test for scale effects Team et al. (2025).
Table A1: Models used in our evaluation, along with their context-window sizes and selection rationale.
Table A2 lists the cut-off for the models used for our hallucination study.
Model Cut-off date (training data freshness)
Llama3-8b-instruct Mar 2023
Llama 3.2-3b-instruct Oct 2024
Phi-3-mini-4k-instruct Oct 2023
Table A2: Models and their cut-off dates for training data freshness.
A.3 Further Information on our NewLensSync Dataset
We use a fixed list of trusted domains: bbc.co.uk, reuters.com, apnews.com, npr.org, pbs.org, theguardian.com, wsj.com, nytimes.com, and propublica.org.
The NewsLensSync orchestrator builds a datastore for each article containing source, author, title, api_description (from NewsAPI), webscraped_description (extracted directly from the article to overcome the API limit), falsified_description (negated version for hallucination testing), api_content, url, and publishedAt.
A scheduled CRON job performs one API call per day over the trusted list, minimizing duplication. Any rare overlaps are removed during preprocessing, retaining only unique records for LLM evaluation. Duplication checking is deferred to evaluation time to avoid overhead. The dataset continues to self-update for future research use.
A.4 Descriptive Statisics on Amazon Reviews and Information on MediaSum News datasets
Dataset Mean Median SD Min Max
Amazon Reviews 2180.3 2185.5 354.5 790 3169
MediaSum News 1166.1 1054.0 621.8 89 2924
Table A3: Descriptive statistics of text lengths (tokens) for Amazon Reviews and MediaSum News.
A.5 Statistical Test Results for Primacy Bias
Model Dataset T-Statistic P-Value
Llama-3-8B-Instruct Amazon Reviews 28.2896 8.62e-130
Llama-3-8B-Instruct MediaSum News 26.8965 2.64e-120
Llama-3.2-3B-Instruct Amazon Reviews 28.3100 6.26e-130
Llama-3.2-3B-Instruct MediaSum News 25.1535 1.50e-108
Phi-3-mini-4k-Instruct Amazon Reviews 18.2547 1.81e-64
Phi-3-mini-4k-Instruct MediaSum News 19.9294 1.14e-74
Qwen3-4B-Instruct Amazon Reviews 24.9284 4.82e-107
Qwen3-4B-Instruct MediaSum News 21.6017 3.18e-85
Gemma-3-27B-IT Amazon Reviews 27.7788 2.64e-126
Gemma-3-27B-IT MediaSum News 24.1413 1.01e-101
Table A4: Paired t-test results comparing the cosine similarity between the summary and the beginning of the source text ($s_{b}$) versus the middle of the source text ($s_{m}$). The statistics shown are for the baseline performance of each model.
To validate that the observed primacy effect (where summary-to-beginning similarity, $s_{b}$, is higher than summary-to-middle similarity, $s_{m}$) is a systematic bias, we perform paired t-tests for each model and dataset combination. Table A4 presents the detailed statistical results. $p\ll.001$ across all tested models confirms that the higher similarity scores for the beginning segment are statistically significant, and the large t-statistics further indicate a substantial effect size.
A.6 Detailed Descriptions of Mitigation Methods
Family Method Operational definition Target bias Hyper-parameters Citation
Prompt Self-Awareness Prompt Prepend an explicit directive that the model remain neutral, preserve sentiment, and cover all parts of the source text. General Mair et al. (2014); Echterhoff et al. (2024)
Chain-of-Thought Prompting Model generates intermediate reasoning steps for each segment before emitting a final summary, encouraging balanced coverage. General Wei et al. (2023)
Cloze-Style Prompt Insert tags BEGIN:__, MIDDLE:__, END:__; model fills blanks, then emits FINAL_SUMMARY. Primacy Tags inside prompt; no extra threshold. Liu and Lapata (2019)
Cognitive Counterfactual Simulation Draft $\rightarrow$ imagine how primacy, recency, or framing bias would distort it $\rightarrow$ rewrite to avoid those distortions. Primacy,   Framing Two passes. Dong et al. (2023)
Self-Help Debias Draft, model self-critiques positional coverage, rewrites summary. Primacy Rewrite pass 300 tokens. Echterhoff et al. (2024)
Task-Specific CoT Prompt Calibration Establishes clear evaluation criteria by explicitly specifying when to return “True” or “False” during fact-checking tasks. For example: “Return true only if the described event has occurred, or if it is a direct consequence of a previously known event”. Return false in all other cases Hallucination Chain-of-Thought sequence pertaining to your task Wei et al. (2023)
Knowledge Boundary Awareness Explicitly defines the model’s knowledge cutoff date to create clear temporal boundaries for what the model can reasonably be expected to know. Prevents models from fabricating information about events beyond their training data by acknowledging these boundaries in prompts. Hallucination Knowledge cutoff date specification for the model being employed
Epistemic Tagging Requires models to express confidence levels alongside factual assertions, creating a more nuanced representation of the model’s knowledge state. Forces the model to evaluate its own fact-checking abilities before responding and provides users with meta-cognitive signals about response reliability. Hallucination Confidence level scales (we used high and low). This can include moderate, very high, very low depending on task. Lin et al. (2022)
Chunk Partial Summaries Ensemble Split the document into equal-length chunks and summarize each chunk independently (same generation hyperparameters). In a second pass, we merge the partial summaries into a final summary. Primacy Chunk size $\lfloor|D|/3\rfloor$; greedy merge. Ou and Lapata (2025)
Weighted Summaries We pre-allocate the summary length budget so that 33% focuses on the beginning, 34% on the middle, and 33% on the end. The model is instructed not to exceed these chunk limits, which helps prevent overemphasis on earlier parts. Primacy Token ratio 0.33:0.34:0.330.33{:}0.34{:}0.33. inspired by Liu and Lapata (2019)
Re-rank Attention-Sort Re-ordering Run a forward pass to estimate cross-attention weights for each paragraph or segment, then reorder segments from lowest-attended to highest-attended. The model is then re-prompted with this new order, which pushes under-attended parts later in the context and encourages more balanced coverage. Primacy 2 iterations; paragraph granularity.
Position-Invariant Shuffle Shuffle entire sentences (split by periods) to randomize their order before prompting the model, so it cannot rely on absolute position. This can disrupt semantic continuity slightly, but ensures that coverage does not depend on text location. Diagnostic ablation for positional sensitivity. Primacy Period-split; seed $42$. Wan et al. (2024); Liu et al. (2024)
Decode Forced Balanced Coverage We measure coverage of the beginning, middle, and end sections via TF-IDF Salton and Buckley (1988), and add $\log\gamma$ ($\gamma=1.5$) to logits of tokens from under-covered sections until $\lvert s_{b}\!-\!s_{e}\rvert\leq 0.05$. Primacy Threshold $0.05$; boost $\gamma=1.5$. See et al. (2017)
Weighted Token Decoding Multiply next-token probabilities by weights $w_{i}$ (down-weight negative words, up-weight middle keywords). Framing $w_{\text{neg}}=0.3$, $w_{\text{mid}}=2.0$. Liu et al. (2021); Dathathri et al. (2020)
Mirostat Decoding After each token, compute its surprise $s_{t}=-\log p_{t}$; update the running state with $\mu_{t+1}=\mu_{t}-\eta\,(s_{t}-\mu^{\star})$, set temperature $T_{t}=\exp(\mu_{t+1})$, and rescale logits as $\mathrm{softmax}(z/T_{t})$. The feedback loop keeps the observed surprise near the target $\mu^{\star}$, smoothing positional coverage. Primacy $\mu^{\star}=2.0,\;\eta=0.1$ Basu et al. (2021)
Rejection Sampling If top-1 token would increase chunk imbalance, set its logit to -$\infty$ and resample within top-$k$. Primacy $k=5$.
Self-Debias Decoding Dual-pass decoding: at each step run a second forward pass on the current context preceded by a bias-inducing prefix that names the undesired attribute; obtain $p_{\text{bias}}$, compute $\Delta=p_{\text{main}}-p_{\text{bias}}$, and down-scale tokens with $\Delta<0$ via $\alpha(\Delta)=e^{\lambda\Delta}$. Framing $\lambda=10$; bias prefix <30<\!30 tokens; refresh every 4 steps. Schick et al. (2021)
Local-Explanation Guard After each token the model explains its choice; if explanation says “ignoring middle” or “flipping sentiment”, reject token and resample. Primacy,   Framing Check every 5 steps; rule-based detector.
Table A5: Overview of the 18 bias-mitigation strategies in this work. Methods are grouped family, annotated with the bias they aim to address, and accompanied by main hyper-parameters and primary citation(s).
A.7 Full Mitigation Results
Amazon Reviews MediaSum News
Method $\varphi_{\mathrm{frame}}\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\rho_{\text{pri}}\downarrow$ $\varphi_{\mathrm{frame}}\downarrow$ $\bar{s}_{b}\uparrow$ $\bar{s}_{m}\uparrow$ $\bar{s}_{e}\uparrow$ $\psi_{\mathrm{pri}}\downarrow$ $\rho_{\text{pri}}\downarrow$
GPT-3.5-turbo 16.0% 0.848 0.826 0.825 7.6% 24.8% 0.840 0.823 0.820 6.1%
Llama-3.2-3B-Instruct 14.9% 0.860 0.842 0.840 7.4% 22.1% 0.851 0.837 0.832 5.1%
Self-Awareness Prompt 24.0% 0.908 0.883 0.881 19.7% 80.4% 20.9% 0.907 0.883 0.875 15.7% 85.0%
Chain-of-Thought 23.6% 0.909 0.883 0.881 19.9% 81.1% 22.6% 0.907 0.883 0.876 16.9% 85.9%
Cloze-Style Prompt 23.8% 0.908 0.883 0.881 19.8% 81.0% 22.6% 0.907 0.883 0.876 16.9% 85.9%
Cognitive Counterfactual Sim. 23.7% 0.908 0.883 0.881 20.2% 80.8% 22.5% 0.907 0.884 0.876 16.6% 84.8%
Self-Help Debias 23.0% 0.908 0.883 0.881 20.6% 81.3% 23.6% 0.908 0.884 0.876 15.9% 85.8%
Partial Summaries Ensemble 21.0% 0.886 0.870 0.869 14.8% 69.1% 27.6% 0.871 0.854 0.846 14.3% 78.2%
Weighted Summaries 20.7% 0.921 0.903 0.899 15.4% 75.6% 29.4% 0.876 0.862 0.852 9.8% 80.1%
Attention-Sort Re-ordering 19.8% 0.905 0.886 0.885 16.8% 71.4% 26.4% 0.859 0.841 0.830 14.6% 78.5%
Position-Invariant Shuffle 26.6% 0.903 0.890 0.889 10.8% 67.1% 30.8% 0.875 0.868 0.859 8.9% 68.4%
Forced Balanced Coverage 21.0% 0.905 0.884 0.883 17.5% 72.5% 26.2% 0.891 0.867 0.859 19.0% 83.0%
Weighted Token Decoding 18.9% 0.865 0.846 0.844 17.3% 71.0% 29.6% 0.855 0.841 0.834 14.4% 72.3%
Mirostat Decoding 20.0% 0.918 0.903 0.901 12.2% 70.0% 31.2% 0.901 0.882 0.875 11.7% 79.1%
Rejection Sampling 19.8% 0.904 0.885 0.883 16.7% 72.0% 23.8% 0.890 0.865 0.857 19.5% 82.7%
Self-Debias Decoding 22.9% 0.895 0.879 0.878 14.3% 68.9% 26.0% 0.879 0.854 0.847 17.6% 83.9%
Local-Explanation Guard 20.2% 0.865 0.844 0.842 17.9% 73.5% 26.5% 0.870 0.847 0.836 20.1% 81.8%
Llama-3-8B-Instruct 14.5% 0.855 0.837 0.836 7.0% 21.9% 0.847 0.834 0.828 4.0%
Self-Awareness Prompt 15.2% 0.909 0.881 0.878 22.9% 83.2% 21.0% 0.889 0.864 0.855 19.7% 85.5%
Chain-of-Thought 15.3% 0.909 0.881 0.878 22.6% 82.7% 22.5% 0.890 0.865 0.856 19.2% 85.6%
Cloze-Style Prompt 15.2% 0.909 0.881 0.878 21.9% 82.4% 21.9% 0.891 0.866 0.856 20.2% 85.3%
Cognitive Counterfactual Sim. 14.7% 0.909 0.881 0.878 22.5% 82.7% 21.8% 0.890 0.866 0.856 19.0% 84.8%
Self-Help Debias 15.2% 0.909 0.881 0.879 21.8% 83.0% 21.7% 0.891 0.865 0.856 19.2% 85.4%
Partial Summaries Ensemble 16.5% 0.890 0.871 0.870 15.4% 72.7% 28.5% 0.874 0.858 0.849 14.1% 79.3%
Weighted Summaries 15.1% 0.925 0.904 0.902 15.1% 77.0% 26.6% 0.880 0.866 0.855 9.0% 81.8%
Attention-Sort Re-ordering 16.9% 0.879 0.858 0.857 17.5% 74.7% 25.0% 0.869 0.845 0.831 20.7% 85.0%
Position-Invariant Shuffle 21.2% 0.899 0.888 0.887 11.5% 64.4% 27.0% 0.861 0.854 0.847 8.0% 66.3%
Forced Balanced Coverage 15.6% 0.900 0.878 0.877 19.5% 74.2% 23.8% 0.862 0.839 0.832 18.1% 81.3%
Weighted Token Decoding 13.6% 0.871 0.851 0.852 18.4% 70.0% 24.9% 0.852 0.831 0.823 18.1% 77.3%
Mirostat Decoding 13.8% 0.916 0.896 0.895 15.5% 75.8% 31.3% 0.900 0.875 0.867 18.3% 83.0%
Rejection Sampling 14.3% 0.911 0.890 0.889 18.9% 73.9% 26.8% 0.872 0.849 0.842 18.5% 80.0%
Self-Debias Decoding 16.0% 0.905 0.883 0.882 18.8% 75.1% 25.5% 0.879 0.855 0.848 18.2% 82.8%
Local-Explanation Guard 16.1% 0.859 0.837 0.836 18.9% 74.6% 25.9% 0.852 0.830 0.822 18.3% 79.1%
Phi-3-mini-4k-Instruct 34.5% 0.836 0.822 0.823 5.4% 26.2% 0.828 0.827 0.823 4.9%
Self-Awareness Prompt 25.5% 0.882 0.866 0.867 14.5% 69.5% 24.9% 0.850 0.832 0.826 16.6% 73.7%
Chain-of-Thought 23.9% 0.884 0.868 0.868 14.1% 69.7% 26.0% 0.851 0.832 0.827 15.7% 73.5%
Cloze-Style Prompt 24.7% 0.883 0.868 0.868 13.6% 67.7% 25.9% 0.850 0.832 0.826 16.4% 73.5%
Cognitive Counterfactual Sim. 22.8% 0.877 0.862 0.862 14.4% 67.7% 25.5% 0.850 0.832 0.826 16.2% 73.5%
Self-Help Debias 22.8% 0.879 0.863 0.864 14.3% 68.5% 25.1% 0.851 0.832 0.826 16.0% 75.5%
Partial Summaries Ensemble 28.0% 0.893 0.880 0.875 11.5% 71.3% 27.7% 0.858 0.845 0.836 11.4% 73.9%
Weighted Summaries 22.6% 0.920 0.904 0.899 11.2% 78.1% 27.8% 0.880 0.871 0.863 7.4% 72.1%
Attention-Sort Re-ordering 33.1% 0.887 0.877 0.876 9.9% 66.8% 28.1% 0.837 0.823 0.818 14.5% 67.9%
Position-Invariant Shuffle 28.2% 0.877 0.866 0.865 9.3% 67.6% 28.2% 0.837 0.834 0.825 7.7% 63.0%
Forced Balanced Coverage 31.5% 0.901 0.891 0.893 7.6% 61.5% 29.9% 0.844 0.831 0.830 11.8% 66.3%
Weighted Token Decoding 25.0% 0.860 0.849 0.854 9.8% 56.9% 28.9% 0.841 0.829 0.828 14.4% 63.0%
Mirostat Decoding 33.2% 0.891 0.883 0.884 4.2% 58.8% 34.3% 0.854 0.845 0.843 8.7% 61.3%
Rejection Sampling 22.7% 0.885 0.876 0.878 8.2% 60.5% 28.3% 0.853 0.836 0.835 14.6% 66.8%
Self-Debias Decoding 28.8% 0.894 0.885 0.888 7.1% 56.4% 31.3% 0.844 0.832 0.834 11.3% 60.0%
Local-Explanation Guard 22.5% 0.886 0.875 0.879 9.4% 59.7% 27.3% 0.841 0.828 0.826 12.4% 65.8%
Table A6: Coverage, framing-change, and primacy-bias metrics for Amazon Reviews and MediaSum on $\varphi_{\mathrm{frame}}$ (framing-change fraction, $\downarrow$), $\bar{s}_{b},\bar{s}_{m},\bar{s}_{e}$ (cosine similarity with the first/middle/last third of the source, $\uparrow$), $\psi_{\mathrm{pri}}$ (share of summaries whose similarity to the beginning exceeds that to the middle by $5\%$, $\downarrow$), and $\rho_{\text{pri}}$ (share of summaries exhibiting primacy bias, $\downarrow$).
A.8 Prompt Templates
Prompt Strategy Prompt
Summarization and General Bias Mitigation Prompts
Self-Awareness Prompting “You are an unbiased summarizer. Be mindful not to introduce any framing bias or omit the middle. Preserve the original sentiment. Please summarize the following text: [DOCUMENT_TEXT]
FINAL_SUMMARY:”
Chain-of-Thought (Summarization) “Please read the text below carefully. Then break down the text into beginning, middle, and end, describing each portion in detail. After that, produce a final summary. Use the following format:
BEGIN_ANALYSIS: [describe the beginning]
MIDDLE_ANALYSIS: [describe the middle]
END_ANALYSIS: [describe the end]
FINAL_SUMMARY: [your final concise summary]
Text: [DOCUMENT_TEXT]
Cloze-Style Prompt “Fill the blanks for each part of the text:
BEGIN: ____ MIDDLE: ____ END: ____
Text: [DOCUMENT_TEXT]
When you fill them in, finally produce:
FINAL_SUMMARY:”
Cognitive Counterfactual Simulation (Final Revision) “Original text: [DOCUMENT_TEXT]
Initial Draft: [DRAFT_SUMMARY]
Here are the potential biased deviations: [LIST_OF_SIMULATED_BIAS_DEVIATIONS]
Please produce a final summary that avoids these biases.
FINAL_SUMMARY:”
Self-Help Debias (Rewrite Stage) “Rewrite the following summary to remove any bias, preserve the original sentiment, and ensure equal coverage of beginning, middle, and end: [DRAFT_SUMMARY]
FINAL_SUMMARY:”
Partial Summaries Ensemble (Merge Prompt) “Combine the following partial summaries into one coherent summary:
[PARTIAL_SUMMARY_1] [PARTIAL_SUMMARY_2]
FINAL_SUMMARY:”
Weighted Summaries (Chunk Prompt) “Summarize this portion in about [PORTION_TOKEN_BUDGET] tokens:
[CHUNK_TEXT]
FINAL_SUMMARY:”
Attention-Sort Re-ordering (Final Prompt) “CONTEXT:
Segment 1: [SORTED_SEGMENT_1_TEXT]
Segment 2: [SORTED_SEGMENT_2_TEXT]
TASK: Summarize:
FINAL_SUMMARY:”
Position-Invariant Shuffle (Prompt) “The text is out of order; please summarize it fully: [SHUFFLED_DOCUMENT_TEXT]
FINAL_SUMMARY:”
Hallucination Mitigation Prompts
Baseline “You are a fact-checking AI. For each statement, predict if the event has occured or not and respond with ONLY ’True’ if it’s factual or ’False’ if it’s not. Do not provide any explanation. Respond with TRUE or FALSE only and nothing else.”
Prompt Calibration (CoT) “You are a fact-checking AI. Follow these strict guidelines:
- Respond with ’True’ ONLY if:
* You are 100% certain the statement describes an event that has occurred
* You have clear knowledge of the event from your training data
* You can identify specific details confirming the event happened
- Respond with ’False’ if:
* You have ANY doubt about the statement’s accuracy
* The statement contains ANY misleading or incorrect information
* You lack reliable information about the statement
Respond with EXACTLY one word: ’True’ or ’False’. No explanations.”
Knowledge Boundary Awareness “You are a fact-checking AI with knowledge up to {knowledge_cutoff}. For each statement, predict if the event has occurred or not. We check your decision-making abilities on futuristic events based on current data. Respond with only ’True’ or ’False’ for each news event.
If the event is scheduled to occur after {knowledge_cutoff}, make a good prediction based on your knowledge up to {knowledge_cutoff}.
RESPOND WITH EXACTLY ONE WORD: ’True’ or ’False’.”
Epistemic Tagging “You are a fact-checking AI. Analyze this statement and respond with:
- ’True [High Confidence]’ or ’True [Low Confidence]’
- ’False [High Confidence]’ or ’False [Low Confidence]’
Use these epistemic tags to guide your reasoning:
[Certain] = Information you know with high confidence based on your training data
[Uncertain] = Information you’re less confident about or may be outside your knowledge
[Reasoning] = Step by step analysis to determine veracity
[Conclusion] = Your final determination with confidence level
Respond ONLY with one of the four exactly formatted options listed above and nothing else.”
Table A7: Prompt strategies and the prompts used in our experiments.
A.9 Experimental Details on Products for Human Study
Table A8: Product reviews and their summaries used for the human-subject experiment.
Product Review Summary
Refrigerator water filter We installed this filter on our Whirlpool Refrigerator and two months later the water was only trickling. I reseated it just in case something had gotten in the line. Nothing had gotten in. The filter was plugged after two months. The valve and all are fine. After reseating water was still trickling. We ordered and installed the factory Whirlpool filter and as expected the water flow is normal. The waster valve made noise in the days that this cheap filter was clogged, think of the damage that does over time! Why take a chance with this cheaper unit coming apart inside and risking a Service Call because it damaged the valve that is hard to get to?????<br /><br />PAY TYHE MONEY. YOU BOUGHT A NICE FRIDGE. BUY THE WHIRLPOOL FACTORY FILTER. IT IS YOUR HEALTH, IF THIS FILTER CLOGGS AFTER TWO MONTHS THEN IT IS FAILING INSIDE. The old whirlpool filter went over one year not changed and never had an issue. I now have the Whirlpool filter in my subscribe-n-save every six months. I initially tried a cheaper filter for my Whirlpool Refrigerator, but after two months it clogged and caused water flow issues. I switched to the factory Whirlpool filter and the water flow returned to normal. It’s worth investing in the quality filter to avoid potential damage and service calls.
Coffee water filter Three stars ONLY because the filter assembly was almost impossible for me to get open. If you can’t replace the filter, the filter assembly is useless. The bottom of the filter assembly requires the sides of the hard acrylic be squeezed together in order for the bottom to be released (refer to the attached photo). After several failed attempts, I was able to get the bottom released by placing it against a hard object and pressing down on the opposite side with both thumbs. Very worrisome because the acrylic could just as easily cracked, ruining the filter assembly. So, the bottom line is if you don’t have strong hands, it will be very hard to open the filter assembly to replace the filter every two months. Otherwise, this starter kit fits nicely in my Keurig 2.0, K300, the water is clean and tastes very good. The reusable K-cup is another plus for the starter kit. My Keurig recognizes it and allows for brewing loose leaf tea or non-K-cup coffee of my choice. The filter assembly was difficult to open, requiring strong hands to avoid cracking the acrylic. However, the water tastes great and the reusable K-cup is a nice addition for brewing loose leaf tea or non-K-cup coffee. Fits well in my Keurig 2.0, K300.
Electric kettle Cons:<br /> * The first thing you will notice on unpacking is the short cord. It is the shortest appliance cord I have ever seen and is quite inconvenient.<br /> * The build of the kettle is not great, especially considering the price. I am worried that the gooseneck will fall off.<br /> * The electronic controls are clumsy when compared to the Cuisinart Perfect Temp simple presets.<br /> * The heating element is small so heating is slower than expected.<br /> * Bonavita designs never get it quite right. I use their coffeemaker which works great but has a smoked glass water reservoir that makes it hard to get an exact water level. The waterproof cover for the electronics on this kettle are frosted instead of clear so they are not easy to see.<br /><br />Pros:<br /> * Heats water to very close to the desired temps.<br /> * Gooseneck is great for pourover coffee preparation.<br /> * Takes up little counter space but must be very close to an outlet. The short cord and build quality are disappointing, but the Bonavita kettle heats water well and is compact. The gooseneck is perfect for pour-over coffee. Controls are not as user-friendly as other brands. Overall, it’s a decent option for those who prioritize temperature accuracy and pour control.
Screen protector Not overwhelmed nor am I disappointed. This protector is on par with many on the market. Installation is all about preparation. The included guide frame is nice and takes guesswork out of alignment. If you’ve prepped the screen and manage to not stir up dust then you should get an installation that only needs bubbles squeezed out. Take your time and you can get a perfect installation. Take your time.<br /><br />That being said, the frame is all that really sets this one apart from others for this iPad and the others we own. You will get the same quality as if you pick another brand as they’re all pretty much the same one in different packaging.<br /><br />It works and looks as it should. After a bit of use you’ll see fingerprints start to show up. It will happen. You can minimize this by using a clean soft cloth with no cleaner. Eventually though it will start to show every print. Up to you to live with it or install the second one. This screen protector is decent, with easy installation using the guide frame. Quality is similar to other brands. Fingerprints show up after use, but can be minimized with a clean cloth. Overall, it works as expected but may require a second installation for a flawless look.
Dryer vent hood The Good:<br />The construction is excellent. Thick stainless steel and well made.<br />The damper works well. We are using this as a dryer vent and the damper moves out fully with even the mildest air flow. Its great not to have a screen that clogs with lint since this is installed on the second floor and getting to the old vent was a big hassle.<br /><br />The downsides:<br />1. The "flapper" oval foam seal that reduces noise and, more importantly, seals the vent when closed came off immediately. With the first use, it went flying off. The important thing is that without this foam, there is a decent sized gap at the top of the flapper that bugs and outside air could get into. I reattached it with some silicon and hopefully that does the trick. The manufacturer needs to use a better adhesive.<br />2. The rubber seal at the base makes installation a challenge. It is installed with friction and comes off with the slightest bump. I had to make quite an effort to keep it in place during install. Note: I wish I had glued it on the day before, so it stayed in place during install rather than fighting it. The stainless steel construction is top-notch and the damper works perfectly for dryer vent use. However, the foam seal came off easily and the rubber seal at the base made installation difficult. Overall, a great product with minor flaws that can be easily fixed with some DIY solutions.
Video doorbell I’ve had a lot of frustration with this product and was considering returning it. It required a lot of troubleshooting after installation based on what was within its view. I suspect the original firmware had problems with audio/video quality that was fixed with a factory reset and registration as a new doorbell. A WiFi router sets about five feet behind the doorbell. Motion detection was triggered at sunrise every day, waking my dog and I up, and I was faced with a puppy who wanted to interact with me an hour before my alarm went off. This was fixed a few weeks later by installing a wedge kit to angle the camera so a wall was not in view.<br /><br />Pros - Easy setup. Get notified by phone. Easy to download video of events. Able to talk with people. Can setup active motion zones. Night vision. Fish eye lens covers a large area. Can snooze motion detection while you cut grass or hang out on the porch. Can set schedule of when device operates. Friendly staff via email/twitter.<br /><br />Cons - Initial problems with audio/video quality. Walls trigger motion detection at sunrise regardless of motion zones setup within the app. Limited choice of sounds on device and smartphones. I initially had issues with this product but after troubleshooting and adjusting settings, it improved. Easy setup, notifications, video download, two-way talk, motion zones, night vision, wide coverage, snooze feature, scheduling, and helpful customer service. However, initial audio/video problems and motion detection triggered by walls at sunrise were frustrating. Limited sound options.
Radio Overall it’s a well working radio. Sound is decent and clear. Display is nice looking. However the only knob is a multi function nightmare. Gone are the days of just reaching over to adjust squelch or volume with a quick turn of a knob. You got to choose the function then turn knob. Really irritating at least. The function knob is also very stiff and doesn’t turn smoothly. This radio is getting pulled out of main driver and thrown into my pickup. I really don’t like it. A shame Anytone was unable to give you two knobs, Volume and Squelch. Recommendation: Spend an extra hundred and buy a better radio.<br />Follow-up: 9/2/18 - I decided to give it a second chance. I pulled this radio off the shelf and immersed myself into programming some frequencies by hand. Took some time figuring out the quirks, like you need to program UHF frequencies when your on B radio. If you try to do it on A, your split will be 600 mhz instead of 5 khz. Well after all this learning I decided to mount it in my car. Its really not to bad once its programmed. I would however recommend a larger external speaker though since the small size radio just cant put out good sound. The other irritating thing is Channel names are limited to 5 characters. Yikes. Initially disappointed with the single function knob, but after giving it a second chance and learning its quirks, I found it to be a decent radio once programmed. Sound quality is lacking, but overall not bad for the price. Would recommend a larger external speaker for better sound.
Reusable K-Cup This isn’t compatible with some machines, out of the box. Keurig has been increasing the size of the top needle, or possibly moving it slightly off-center on some models. When using a K-Cup, that never matters of course. When using a reusable K-Cup like this one, the machine’s lid might not close without ruining the machine … or without wearing down either the needle or the reusable K-Cup’s lid opening.<br /><br />They’re all copying each other’s designs, because of course they are. Who’s gonna be brave, and the first to just make the top opening a radical 1 mm wider?<br /><br />The "smoosh down" method of securing the lid isn’t the best, however I’m aware they’re all like that. If using a disposable filter, push it down into the cup, otherwise the lid will not stay shut. Accept maybe a little bit of sediment in your coffee mug as a result. Yeah, I know, that’s what a disposable paper filter is supposed to mitigate. This reusable K-Cup may not fit all machines perfectly, but works fine with K-Cups. The design is similar to others, but could benefit from a wider top opening. The lid may need extra pressure to close, especially with a disposable filter. Overall, it gets the job done with minor inconveniences.
Headlamp I got this headlamp mainly for trail running in the dark. It’s a bright light that has more of a surround light. There are different settings for brightness. There is also the ability to turn on just the center lights which is better for focusing on what is in front of me on the trail. These are the pros on the light.<br /><br />Unfortunately, there are cons that are big enough that made me drop this light a couple of stars. First of all, it’s stiff on my head and not really comfortable for trail running or hiking. The battery pack sits kind of heavy on the back of my head. I’m not a fan at all of the charging system with the prongs that need to line up just right. It would have been so much easier to make this a regular USB charging system that would have been compatible with other charging cords.<br /><br />This will be going in my camping trailer and used for camping. On the plus side, this will be one of the brightest headlamps I’ve used for camping and will make a great addition to my camp supplies. This would also work really well for those working outside in the dark when brightness is more important than comfort. This headlamp is bright with different settings, great for trail running. However, it’s uncomfortable and has a finicky charging system. Despite the cons, it’s perfect for camping and outdoor work where brightness is key.
Headset I have the Jabra 35 headset also and this one is noticably different with regard to the quality of phone calls, but that’s really the only material difference I can find in my day to day use. The other unit does offer some noise cancellation effects that this one doesn’t have and people you call can definitely tell the difference. My husband said that he could hear background noise much more clearly when I called him from this 15 than with the 35.<br /><br />As far as battery life, this thing is really solid. I’ve used for three hours on two different days without recharging between days. I primarily use it to listen to audiobooks while I’m doing my barn chores and checking fence lines. I have noticed that I’m pretty limited with the distance I can stray from my physical phone before this headset begins to stutter to lose signal - it seems to be about 20 feet with my Sony Xperia phone. The Jabra 15 headset has great call quality but lacks noise cancellation. Battery life is impressive, lasting three hours over two days. However, signal range is limited to about 20 feet from my phone. Overall, a solid choice for calls and audiobooks, but lacking in some features compared to the 35.